### v1

Abstract
Context. The search of exoplanets around nearby M dwarfs is a crucial milestone to perform the census of planetary systems in the vicinity of our Solar System.
Aims. Since 2018 our team is carrying a blind search program for planets around nearby M dwarfs with the near-IR spectro-polarimeter and velocimeter SPIRou at the CFHT and the optical velocimeter SOPHIE at the OHP in France. The aim of this paper is to present our results on Gl 410, a 0.55 M⊙ 480150 Myr old active M dwarf distant 12 pc.
Methods. We search for planetary companions using radial velocities (RVs). We use the line-by-line (LBL) technique to measure the RVs with SPIRou and the template matching method with SOPHIE. Three different methods, two based in principal component analysis (PCA), are used to clean the SPIRou RVs for systematics. Gaussian processes (GP) modeling is applied to correct the SOPHIE RVs for stellar activity. The and apodize sine periodogram (ASP) analysis is used to search for planetary signals in the SPIRou data taking into account activity indicators. We analyse TESS data and search for planetary transits.
Results. We report the detection of a sin=8.41.3 M⊕ sub-Neptune planet at a period of 6.0200.004 days in circular orbit with SPIRou. The same signal, although with lower significance, is also retrieved in the SOPHIE RV data after correction for activity using a GP trained on SPIRou’s longitudinal magnetic field () measurements. TESS data indicates that the planet is not transiting. We find within the SPIRou wPCA RVs tentative evidence for two additional planetary signals at 2.99 and 18.7 days.
Conclusions. Infrared RVs are a powerful method to detect extrasolar planets around active M dwarfs. Care should be taken however to correct/filter systematics generated by residuals of the telluric correction or small structures in the detector plane. The LBL technique combined with PCA offers a promising way to reach this objective. Further monitoring of Gl 410 is necessary.
Key Words.:
Stars: planetary systems; Stars: low-mass; Methods: observational; Techniques: spectroscopic, radial velocities.1 Introduction
One of the most interesting open questions in current Astronomy is the census of extrasolar planets in the Solar System vicinity. The identification of closest extrasolar planetary systems is a crucial milestone for the study of exoplanet atmospheres and biomarkers using future instrumentation combining high-dispersion spectroscopy with high contrast imaging (e.g., Kasper et al., 2021; Chauvin, 2024). As the most numerous stars in the Galaxy, thus in Solar neighbourhood, are low-mass stars ( M⊙), in recent years, there has been a growing interest in extending the planet search effort from Solar-like stars to M dwarfs.
Radial velocity (Bonfils et al., 2013; Pinamonti et al., 2022, Mignon et al. 2024 sub.) and transit surveys (Dressing & Charbonneau, 2013; Mulders et al., 2015; Dressing & Charbonneau, 2015; Hsu et al., 2020; Ment & Charbonneau, 2023) have revealed that a large fraction of M dwarfs, if not all, host a planetary system, and have provided strong evidence supporting the hypothesis that the rocky planet ratio per star is indeed larger than one. Extrapolating the results from field stars to our stellar vicinity is extremely encouraging, as it suggests that is highly likely that a large fraction of nearby M dwarfs host multi-planetary systems. Nearby stars are distributed all over the sky at random orientations. A small fraction of sources are expected to present planets in transit. Missions such as TESS and in the future PLATO will detect a significant fraction of those transiting planets. Nevertheless, to establish the planetary systems census of our closest stellar neighbours, techniques such as astrometry or radial velocities (RV) need to be used in addition to transits.
Surveys for planetary systems around nearby bright M dwarfs with high-precision velocimeters have been carried so far mainly in the optical domain. Recently the RV search has been extended to the near infrared (IR) domain, with the start of operations of new velocimeters such as HPF (Mahadevan et al., 2014), CARMENES (Quirrenbach et al., 2018), SPIRou (Donati et al., 2020) and NIRPS (Bouchy et al., 2017). High-precision IR RVs enables us to explore lower mass stars, sometimes inaccessible for precise RV work in the optical due to their faintness. Furthermore, IR RVs have the potential of being less affected by stellar activity than optical RVs specifically for highly or moderately active stars (e.g. Carmona et al., 2023).
With more than two-thousand low-mass M dwarfs in the 25 pc vicinity, the effort to monitor all these objects is colossal. Thanks to large programmes with the near-IR spectra-polarimeter and velocimeter SPIRou at CHFT (the SPIRou Legacy Survey, the SPICE large programme) and a series of PI SPIRou follow-up programs our team has been monitoring a selected sub-sample of around 60 nearby M dwarfs. At the time of writing, we have obtained more than 100 visits per source for around 45 of them.
In this paper, we report the detection of a sub-Neptune mass planet in a 6 day period around the nearby (12 pc) 0.55 M⊙ 480 Myr old M dwarf Gl 410 with SPIRou, and the tentative detection of two additional planetary signals at 2.99 and 18.7 days. We complement our near-IR data set with quasi simultaneous optical observations obtained with the velocimeter SOPHIE. The paper is organised as follows. We start describing the physical properties of Gl 410 in Sect. 2; then, in Sect. 3, we present the details of the optical and infrared RV measurements. In Sect. 4, we analyze our RV measurements and test the robustness of the planet detection. Then in Sect. 5, we search for the 6 day period signal detected with SPIRou in the SOPHIE data using Gaussian Processes activity filtering techniques. We derive the orbital and physical properties of the detected 6.02 day planet using MCMC techniques in Sect. 6. In Sect. 7, we investigate the presence of other planetary signals within the SPIRou wPCA RVs. In Sect. 8, we present TESS observations of Gl 410 and investigate whether Gl 410b is transiting. We discuss our findings in Sect. 9 and finally, in Sect. 10, we provide our conclusions.
2 Gl410
We summarize the basic physical properties of Gl 410 (DS Leo, HD 95650) in Table 1. Gl 410 has a spectral type M1.0V and a mass of 0.550.02 M⊙ (Cristofari et al., 2022). It is a nearby star located at a distance of 11.9327 pc (Gaia DR3, Bailer-Jones et al., 2021; Gaia Collaboration et al., 2023). Gl 410 is known to exhibit cyclic photometric variability in the optical with periods of about 14 d which is interpreted as caused by rotational modulation of starspots (Fekel & Henry, 2000). The rotational period of Gl 410 has been tightly constrained using spectropolarimetry observations in the optical with ESPaDOnS, NARVAL and HARPS (Donati et al., 2008; Hébrard et al., 2016), and the near-IR with SPIRou (Fouqué et al., 2023; Donati et al., 2023). The latest estimation of is of 13.910.09 days (Donati et al., 2023). Gl 410 displays surface differential rotation with periods at the equator and pole of 13.370.86 and 14.961.25 days respectively (Hébrard et al., 2016). Based on gyrochronogly Fouqué et al. (2023) derived an age of 0.890.1 Gyr for Gl 410. However, there is the possibility that Gl 410 is in fact younger. In Sect. A in the Appendix, we provide further details. We find that empirical rotation-temperature sequences in clusters of known ages indicate that Gl 410 is 480150 Myr, which ranks it among the youngest stars in the solar neighbourhood (Silva Aguirre et al., 2018; Chen et al., 2020), and suggest that Gl 410 might be related to the Ursa Major (UMA) cluster which is approximately 400 Myr old.
Cristofari et al. (2023) based on SPIRou observations measured a average magnetic small scale field strength ¡¿ of 0.71 0.03 kG for Gl 410. Together with the M dwarfs GJ 1289 and GJ 1286, Gl 410 is part of most magnetic stars among the M dwarfs studied in the ¡¿ surveys of Reiners et al. (2022) and Cristofari et al. (2023). The magnetic activity of Gl 410 is further indicated by the presence of emission in the optical spectra. Gl 410 is part of the top one third of the sample in terms of emission strength among the 331 M-stars surveyed by Schöfer et al. (2019).
Zeeman Doppler Imaging (ZDI) investigations by Donati et al. (2008), Hébrard et al. (2016) and Bellotti et al. (2024) show that the large-scale magnetic field of Gl 410 rapidly evolves both in strength and in topology. Between 2007 and 2008, the field had a predominantly toroidal geometry with an average strength of 100 G (Donati et al., 2008). In 2014, the field remained mostly-toroidal with axisymmetric geometry but the average field strength decreased to 60 G (Hébrard et al., 2016). In 2020b to 2021a the magnetic energy became almost equally distributed between the poloidal and the toroidal components and the average field strength decreased to 44 G (Bellotti et al., 2024). In the latest measurements from 2021b to 2022a, the poloidal component became much stronger than the toroidal component with 73% of the magnetic energy and the average field strength further decreased to 18 G (Bellotti et al., 2024).
| Parameter | Value | Ref. |
| RA (J2000) | 11:02:38.34 | 1, 2 |
| DEC (J2000) | +21 58 01.70 | 1, 2 |
| Proper motion [mas] | 142.771 -51.661 | 1, 2 |
| Parallaxes | 83.76420.0204 | 1, 2 |
| Distance [pc] | 11.9327 | 3, 2 |
| Effective temperature [K] | 384231 | 4 |
| log | 4.870.05 | 4 |
| [M/H] | 0.05 0.10 | 4 |
| [/H] | 0.05 0.04 | 4 |
| Mass [M⊙] | 0.550.02 | 4 |
| Radius [R⊙] | 0.5430.009 | 4 |
| Luminosity [ log ] | -1.2390.002 | 4 |
| Rotational period† [days] | 13.870.08 | 5 |
| 13.910.09 | 6 | |
| 13.930.09 | this work | |
| Age [Myr] | 890120 | 5 |
| 480150 | this work | |
| Brightness [mag] | 1, 7 | |
| 1, 7 | ||
| 1, 8 | ||
| 1, 8 | ||
| 1, 8 |
3 Observations
3.1 SOPHIE
Gl 410 was observed during two periods 9 years apart with the high-resolution optical velocimeter SOPHIE222Spectrographe pour l’Observation des Phénomènes des Intérieurs stellaires et des Exoplanètes, meaning, spectrograph for the observation of the phenomena of the stellar interiors and of the exoplanets. (Perruchot et al., 2008; Bouchy et al., 2013) at the 1.93m Telescope of the Haute Provence Observatory in the South of France. The first period with 45 visits was observed from March 14, 2010, to June 14, 2012. A second observation period was planned to coincide with the SPIRou observations, with 58 visits performed from January 26, 2021, to March 04, 2023. Spectra were reduced using the SOPHIE’s data reduction system (Bouchy et al., 2009; Heidari, 2022). RV measurements were obtained using a template-matching technique implemented in the code NAIRA (Astudillo-Defru et al., 2015) which has been adapted for SOPHIE (Hobson et al., 2018). We provide a summary of the SOPHIE the RV measurements in the Table 10 in the Appendix. In Table 2, we provide a summary of the properties of the RV time-series. Figure 1 displays the median subtracted SOPHIE RV data.
3.2 SPIRou
Gl 410 was observed with the near-IR velocimeter and polarimeter SPIRou333SPectropolarimètre InfraRouge. (Donati et al., 2020) mounted at the Canada-France-Hawaii Telescope atop Mauna-Kea during 156 visits from April 2018 to July 2023. Data were obtained in the context of the SPIRou’s commissioning in 2018, the SPIRou Legacy Survey from 2019 to 2022A, and a series of PI follow-program programs (PI. Carmona, PI. Artigau) in 2022B, 2023A and 2023B. SPIRou nightly observations typically consist of a Stokes V sequence of four exposures, each taken with different positions of the Fresnel rhombs inside the polarimeter. The combination of the channel A and channel B of the four sub-exposures enables to recover the Stokes V (and Stokes I) spectra for the magnetic field analysis. For each science observation, a simultaneous Fabry-Pérot exposure is obtained in the calibration channel C. In the context of the radial velocity analysis, we use the combined extraction of the science channels A and B and each sub-exposure of the Stokes V sequence is treated independently.
For the radial velocity analysis we used the data reduced using the APERO (Cook et al., 2022) pipeline version v07275 installed at the SPIRou Data Centre hosted at the Laboratoire d’Astrophysique de Marseille. APERO extracts the combined science channel A and channel B spectra, performs the wavelength calibration using nightly observations of the Uranium-Neon lamp and Fabry-Pérot exposures, and do the telluric correction. The telluric correction uses the combination of an atmospheric absorption model generated by TAPAS (Bertaux et al., 2014), a empirical template of the star deduced from observations at different BERV, and a library of nightly observations of telluric standard stars taken since the commissioning of the instrument in 2018. Further details on the telluric correction are given in Artigau et al. (2014), Cook et al. (2022) and Artigau et al. (in prep.). Individual radial-velocity measurements were obtained using the LBL technique (Artigau et al., 2022) and they are corrected by the BERV and the drift of the Fabry-Pérot at each exposure. A further correction for the instrumental radial-velocity zero point is done using a Gaussian Process modeling based on observations of a sample of radial-velocity standard stars observed by SPIRou (Cadieux et al., 2022).
For the analysis of the stellar activity, we used the simultaneous longitudinal magnetic () field measurements obtained with the Libre-ESpRIT pipeline444The nominal ESPaDOnS data reduction pipeline (Donati et al., 1997) adapted for the SPIRou data (Donati et al., 2020). previously published in Donati et al. (2023).
3.3 TESS
The Transit Exoplanet Survey (TESS, Ricker et al., 2014) observed Gl410 (TIC 97488127) with two-minute cadence during Sector 22, from February 18th to March 18th of 2020. We obtained the Presearch Data Conditioning (PDC) flux time-series (Smith et al., 2012; Stumpe et al., 2012, 2014) processed by the TESS Science Processing Operations Center (SPOC, Jenkins et al., 2016) from the Mikulski Archive for Space Telescopes (MAST)555mast.stsci.edu.
| SOPHIE1 | SOPHIE2 | SOPHIE 1+2 | |
| Time-series | |||
| N | 45 | 58 | 103 |
| ¡RV¿ [km s-1] | -13.8399 | -13.8347 | -13.8385 |
| ¡¿ [m s-1] | 2.5 | 2.4 | 2.4 |
| rms [m s-1] | 14.6 | 12.7 | 13.9 |
| LS periodogram | |||
| [days] | 6.99 | 6.91 | 6.92 |
| norm. power | 0.37 | 0.42 | 0.33 |
| FAPbootstrap | 0.2 | 3.810-3 | 4.010-4 |
4 Radial velocity analysis
For the radial velocity analysis we utilise Generalized Lomb Scargle (GLS) periodograms (Zechmeister & Kürster, 2009). To determine the false alarm probability levels (FAP) we used two methods. For peaks with significance lower than 10-5 and the 10%, 1%, 0.1% FAP levels in the periodograms, we used a bootstrap algorithm with 105 iterations. The method consists in generating a time series with dates similar to the real series of measurements, but assigning to each of them a RV measurement chosen randomly from the original data series. We label these FAP levels FAPbootstrap. For peaks with significance higher than 10-5 (i.e. inferior to 1/(number of boostrap iterations)), we used the GLS FAP analytical implementation of the package pyAstronomy (Czesla et al., 2019). We label these FAP levels FAPanalytical.
4.1 SOPHIE radial velocities
We have achieved with SOPHIE an average RV precision per measurement of 2.4 m s-1 and a rms of 13.9 m s-1 in the full time-series. Figure 2 displays the GLS periodograms obtained for the full data set, the 20102012 and the 20212023 periods respectively. Table 2 summarizes the properties of the time-series and the strongest peak in the periodogram.
The full data-set displays a peak at days with a FAPbootstrap of 4.0. This peak corresponds to the /2 harmonic of the stellar rotation period of days (Donati et al., 2023). Besides the 1 day alias there are no other peaks in the full data-set periodogram. No signal is observed at the stellar rotation period. The 20102012 season shows no significant periodicities. A broad peak is seen close to 6.99 days but its significance is low (FAPbootstrap=0.2). In the 20212023 period, the days appears at a FAPbootstrap level of 3.810-3. Besides the 1 day alias, no peaks are present in the periodogram with significance below a FAPbootstrap=10% level.
In summary, in the raw SOPHIE RV data, there is a signal at 6.9 days period but most likely related to the stellar rotation.
4.2 SPIRou radial velocities
We have achieved with SPIRou an average RV precision per measurement of 2.3 m s-1 and a rms of 9.1 m s-1 in the full LBL time-series. We display in the left panels of Fig. 3, the SPIRou LBL RV measurements as a function of time and their GLS periodogram. The periodogram displays a peak at days with a FAPbootstrap level of 2.2%. No peaks are observed at the rotational period of the star nor its harmonics. Our SPIRou LBL RV measurements are listed in Table LABEL:GL410_SPIROU_RVtable in the Appendix. A summary of the properties of the LBL RV time-series is given in Table 6.
Due to the reflex motion of the Earth in its orbit (i.e. the BERV), observations acquired at different moments of the year have different velocity shifts with respect to the center of telluric absorption lines. If a star has an intrinsic average systemic velocity with respect to the Sun, the total velocity of the star at the time of the observations is given by
In the near-IR the telluric absorption lines are abundant, thus the stellar photospheric absorption lines are affected by tellurics. The effect of telluric absorption is in first instance to deform the line profiles. After telluric correction, residuals from atmospheric lines, at a much lower level than original, may remain, and even for a perfect correction, the noise is higher in the affected velocity channels. The moment when the spectra is most affected by tellurics is when the atmospheric absorption is at the core of the stellar photospheric lines, that is when is between 10 and 10 km s-1, i.e. when the lines of chemical species common to the Earth’s atmosphere and the stellar photosphere (e.g. H2O, CO) are on the top of each other, in particular, in regions where the optical depth of the atmosphere is the greatest.
We have color-coded in Fig. 3 the RV measurements based . We display in blue the observations obtained when km s-1 and, in red, the observations taken when km s-1. The telluric correction of APERO is good, the LBL algorithm is efficient filtering outliers, and the zero-point correction in time and BERV space further corrects residuals of telluric lines. However, as can be seen in Fig. 3, at the m s-1 precision level, the systematic noise of tellurics is evident. Fig. 3, shows that a large fraction of the outliers in the RV time-series is in red-color. The effect of telluric lines becomes evident when the RV-time-series is plotted as a function of (see Fig. 3). The systematics due to the telluric absorption can be seen by eye, as an up and down correlated-noise pattern.
| full LBL | —Vtot— 10 km s-1 | LBL wPCA | LBL Wapiti | |
| data-set | LBL | |||
| Time-series | ||||
| visits | 160 | 121 | 157 | 157 |
| ¡RV¿ [km s-1] | 13.898 | 13.897 | 13.807 | - |
| ¡¿ [m s-1] | 2.3 | 2.3 | 2.2 | 2.2 |
| rms [m s-1] | 9.1 | 6.5 | 6.8 | 6.3 |
| GLS periodogram | ||||
| Gl 410b | ||||
| [days] | 6.02 | 6.02 | 6.02 | 6.02 |
| norm. power | 0.14 | 0.34 | 0.22 | 0.23 |
| FAP† | ||||
| candidate signal at 18.7 days | ||||
| [days] | 18.76 | 18.73 | ||
| norm. power | 0.14 | 0.09 | ||
| FAP† | 0.016 | 0.8 | ||
If the RV time-series is filtered of the data-points taken at km s-1 (i.e., the red-points in Fig. 3), the rms of the time-series (see Table 6), improves from 9.1 m s-1 for the full time-series to 6.5 m s-1 for the filtered time-series. If the periodogram is made only with observations taken at km s-1, we not only recover the peak at 6.02 days detected in the full time-series, but its significance increases from an FAPbootstrap=0.022 in the full time-series to FAPanalytical=1.4 in the filtered time-series. The case of Gl 410 dramatically shows the effect that telluric residuals can have adding RV noise thus blurring a planet detection.
While filtering data-points taken at km s-1 is a robust and fast way of finding promising planet candidate signals, besides the obvious fact that it takes out a large number data points (39 measurements), it has the limitation that it can boost the power at certain frequencies because specific and periodic times are missing in the time sampling. The RV systematics and correlated noise generated by telluric residuals and detector based glitches and imperfections can be corrected. The LBL framework produces a 2D data set, a time series for each of the thousands LBL line measurements. In this 2D data-set, data-driven techniques such as principal component analysis (PCA) can be employed to further filter, mitigate and correct systematic noise in the RV measurements.
In the context of the SPIRou collaboration, two PCA-based LBL RV correcting algorithms have been developed: Wapiti (Ould-Elhkim et al., 2023) and a weighted PCA method (wPCA, Artigau et al. in prep). We refer the reader to those papers for details on each methodology. The main difference between these two approaches is that Wapiti performs a PCA in the stellar velocity space, whereas in the second method the PCA is applied in the detector pixel space (inspired on Cretignier et al., 2023). We have re-analyzed the LBL RV data of Gl 410 with both of these methods. In the central and right panels of Fig. 3, we display the RV measurements and GLS periodograms obtained with wPCA and Wapiti. In Table 6, we present the summary of the properties of the RV time-series.
Fig. 3 illustrates how wPCA and Wapiti correct for the systematic noise at km s-1 . The rms of the full time-series decreases from 9.1 m s-1 in the raw LBL to 6.8 m s-1 with wPCA and 6.3 m s-1 with Wapiti. The rms of the points at km s-1 (red dots in Fig. 3) decreases from 14.3 m s-1 in the raw LBL, to a 7.1 m s-1 with wPCA and 8.3 m s-1 with Wapiti. We also observe that the peak at 6.02 days becomes stronger in the full time-series and that it reaches a FAPbootstrap level of 3.5 Furthermore in the wPCA RV data, a new peak at 18.76 days with an FAPbootstrap level of 1.6% becomes visible. A peak at 18.73 days is also present in the Wapiti time-series, however with a negligible significance (FAPbootstrap=80%).
In summary, the SPIRou RV data shows a clear periodic signal at a 6.02 days. The signal is not a harmonic or alias of the rotation period. There is evidence for a second tentative candidate signal at 18.7 days.
4.3 A test of robustness of the 6 day SPIRou signal.
No periodicities are detected at 6.02 nor 18.7 days in the activity indicators time-series of SOPHIE nor SPIRou (details are provided in Sect. C in the Appendix). We performed an additional test of robustness to exclude that the signal at 6 days is due to activity. If a candidate signal is due to a planet (and not activity), it is expected that as more RV data points are accumulated the significance of the detection increases, and this should not depend on the way how the time-sampling of the measurements is performed.
Using the km s-1 SPIRou wPCA time-series, we performed a test that consisted in checking the evolution of the FAP level of the 6 day peak in 100 different randomly sampled time-series. The construction of a RV time-series, starts by selecting randomly 10 measurements, calculating the GLS periodogram and measuring the FAP of the peak at 6 days. Then one randomly selected measurement is included in the time-series, the periodogram recalculated and the FAP of the peak at 6 days measured. The procedure is repeated until the complete number of visits is reached for a time-series. In this way, we constructed 100 independent evolutions of the FAP as a function of the number of measurements. This test implies to calculate around twelve thousand FAPs, to keep the computation time short, we calculated the FAP level using the analytical approximation implemented in pyAstronomy (Czesla et al., 2019). The analytical FAPs are a factor of a few lower than the FAPs calculated using a bootstrap method.
The result of the test is displayed in Fig. 4. We can see that for all the time-series, the FAP level of the 6 day peak decreases as the number of measurements increases. In Fig. 4, we added the plot of the FAP level in the chronological time-series and the median FAP for each bin of number of measurements. The test indicates that the peak at =6 days in the periodogram is physical, and that it does not depend on the sampling of the time-series. This test gives confidence that the 6 day peak in the periodogram is indeed of Keplerian origin.
5 Stellar activity filtering of the SOPHIE RVs and search for the Gl 410b signal.
As opposed to a signal due to activity, a Keplerian reflex motion caused by the presence of a planet is achromatic. There is a hint of the 6 day period signal in the SOPHIE RV data of Gl 410. If the SOPHIE and SPIRou RVs are plotted phase-folded, both data-sets are consistent. However, the SOPHIE RVs have a much larger dispersion (14 m s-1) than the SPIRou RVs (7 m s-1). One interesting question is whether Gl 410b’s RV periodic signal could be retrieved in the SOPHIE data-set if the RV jitter induced by stellar activity is corrected.
We tested two methods to search for the 6 day signal within the SOPHIE RV data. In the first method, the SOPHIE RVs time-series is analyzed with a model which is the sum of a quasi-periodic Gaussian Process (GP) and a Keplerian model. In the second method, we use the measurements of the longitudinal magnetic field () from SPIRou to train a GP that is used to correct the SOPHIE RVs for activity jitter, and then the planet signal is searched in the RV residuals.
5.1 Method 1: SOPHIE GP modeling using the SOPHIE RV data alone.
As the rotational period of Gl 410 is well constrained by spectropolarimetry, the SOPHIE RV data can be analyzed using a GP + Keplerian model. The goal here is to determine what model is more likely between a model without planet (i.e. only a GP) or a model with GP + planet. To perform the analysis we used RadVel (Fulton et al., 2018), and assumed a circular orbit. A Gaussian prior was used for the GP period with center at 13.9 days and =0.1 days corresponding to the stellar rotation period derived from the measurements of the longitudinal magnetic field (Donati et al., 2023). Uniform priors were set for: the planet period (0 to 20 days); the planet semi-amplitude (0 to 20 m s-1); the time of inferior conjunction (2459648.0 10 days); the GP amplitude (0 to 75 m s-1); the GP decay-time (0 to 2000 days); the GP smoothing-length (0 to 2 days); and the uncorrelated noise (0.0 to 10 m s-1).
Table 4 provides a summary of the AIC tests comparing the models. The model with GP model alone (i.e. without planet) is statistically “ruled-out” by RadVel. In Table 13 in the Appendix, we present the posterior distributions obtained for the the model777 The stellar rotation period is well constrained from spectropolarimetry and provides a stringent prior on the cycle length of the GP which is recovered in the posterior distribution. We tested two models with a broader Gaussian distribution prior for the cycle length (13.93.5 and 13.97 days). All models converged to a GP cycle length of 13.840.03 days. . The Keplerian signal at 6.02 days is recovered. This test provides independent evidence for the existence of the 6 days planet inside the SOPHIE RV data. In Sect. 6.3, we will discuss the coherence of the Keplerian signals obtained with SPIRou and SOPHIE.
| Model | RMS | BIC | AICc | AICs | |
| AICc Favored Model | |||||
| , GP, , | 9 | 4.9 | 450 | 435 | 0.0 |
| Ruled Out | |||||
| GP, , | 6 | 23 | 866 | 855 | 420 |
5.2 Method 2: SOPHIE GP modeling using SPIRou’s as activity proxy.
In this method, the search of the signal is done using a two step approach: in step one, a quasi-periodic Gaussian process (GP) model is fit to the SPIRou’s longitudinal magnetic field () time-series, and in step two, the posterior distribution of the hyper parameters (period, decay time and smoothing length) of the GP are used as prior for a quasi-periodic GP model of the SOPHIE RVs (other hyperparamers, ln , ln , priors are taken as uniform). The aim of this method is to filter the SOPHIE RVs for activity using as a proxy of the activity signal and search for the signal at 6 days in the residuals.
We used the SPIRou measurements from Donati et al. (2023). In Sect. D in the Appendix, we describe the the quasi-periodic GP model used. In Fig. 22, also in the Appendix, we display the MCMC corner plots of the GP. We retrieve in our GP model of a stellar rotational period of days which is consistent with the previous determinations of the rotational period by Donati et al. (2023) and Fouqué et al. (2023). We modeled the SOPHIE RVs with the same quasi-periodic GP model and implementation used for . We focused on the SOPHIE RVs from 20212023 as this period has the quasi-simultaneous SPIRou measurements888Bellotti (2023) investigated in detail the SPIRou measurements of Gl 410 using GPs. One of their conclusions was that the decay time differed in the 2020/2021 and 2021/2022 data sets. Using ZDI, Bellotti (2023) found that the surface differential rotation was stronger in 2021/2022 inducing a faster surface variability of activity tracers. We use here a single GP for the complete time-series. Our GP, therefore, represent an ”average” that does not capture these fluctuating activity levels.. The prior and posterior distributions for the RV GP are given in Table 13 in the Appendix. In Fig. 23, also in the Appendix, we show the MCMC corner plots of the RV GP model. Fig. 5 summarizes the model results. The upper panel shows the GP model (in blue) and the SOPHIE RVs (black points). The middle panel displays the GP corrected SOPHIE RVs. The lower panel shows the GLS periodogram of the GP corrected SOPHIE RVs. After correcting for the activity jitter, the RV rms decreases from 12.7 m s-1 to 6.0 m s-1 and the GLS periodogram show a clear and strong peak at days with a FAPbootstrap of 2.4. This period is the same as that obtained in the SPIRou RV measurements.
In Fig. 6, we display the phase-folded SPIRou RV time-series together with the SOPHIE GP corrected RVs, overlaying a Keplerian model fit with m s-1 and days. The data of both instruments are compatible. This provide us, one more time, further arguments to support the idea that the periodic signal at 6 days detected with SPIRou indeed has a Keplerian origin.
We calculated a periodogram for the combined SPIRou (wPCA) and SOPHIE-GP corrected RV time-series. It is shown in Fig. 7. The periodogram of the combined SPIRou + SOPHIE-GP data-set (black lines) is very similar to that of the wPCA SPIRou data alone (green lines). However, the addition of the SOPHIE-GP RVs makes the FAP levels of the periodogram to decrease to lower powers (note the difference between the horizontal grey and green lines in the plot). The significance of the peak at days increases from a FAPanalytical=3.5 in the wPCA SPIRou data alone to a FAPanalytical=1.5 in the combined near-IR + optical data-set.
We find that the properties of the RV GP trained on are different to those of the GP based on the RV data alone. The period of both GPs is the same, as we imposed in the priors the rotational period of the star. The amplitude of both GP is different 21 m s-1 for the GP based on the SOPHIE RVs alone, and 123 m s-1 RV GP trained on . The decay time of both GPs is significantly different, from 602 days from the GP based on the SOPHIE RVs alone, to 608 days from the RV GP trained on . These differences in the GP will produce different amplitudes of the Keplerian signal recovered. In Sect. 6, we will discuss in detail the properties of the Keplerian signal retrieved in the SPIRou data. We will show in Section 6, that recovered in the SOPHIE-RV data-set corrected with the GP using as proxy of activity it is almost the same that the derived from the SPIRou wPCA and Wapiti data-sets. Our analysis suggests that a GP trained on as proxy of activity provides indeed a good correction for the stellar activity RV jitter in the optical.
| Model | RMS | BIC | AICc | AICs | |
| LBL km s-1 | |||||
| AICc Favored Model | |||||
| , , , | 5 | 5.1 | 770 | 751 | 0.0 |
| Somewhat Disfavored | |||||
| , , | 7 | 5.2 | 767 | 753 | 2.1 |
| Ruled Out | |||||
| , | 2 | 6.5 | 1868 | 1863 | 1112 |
| wPCA | |||||
| AICc Favored Model | |||||
| , , , | 7 | 5.8 | 1034 | 1013 | 0.0 |
| Nearly Indistinguishable | |||||
| , , | 5 | 5.9 | 1029 | 1014 | 0.5 |
| Ruled Out | |||||
| , | 2 | 6.8 | 2422 | 2416 | 1402 |
| Wapiti | |||||
| AICc Favored Model | |||||
| , , | 5 | 6.0 | 1033 | 1018 | 0.0 |
| Nearly Indistinguishable | |||||
| , , , | 7 | 5.9 | 1040 | 1019 | 1.0 |
| Ruled Out | |||||
| , | 2 | 6.9 | 2422 | 2416 | 1398 |
| LBL km s-1 | wPCA | Wapiti | |
| (baseline solution) | |||
| Orbital parameters | |||
| [days] | |||
| [m s-1] | |||
| [JD] | |||
| Planet parameters | |||
| [M⊕] | |||
| [au] | |||
| Other parameters | |||
| [m s-1] |
6 Gl 410b physical properties
6.1 Gl 410b physical properties from the SPIRou RVs
We derived the Gl 410b planet candidate properties using RadVel. We calculated RadVel models for each of SPIRou RV data-sets: LBL km s-1, wPCA and Wapiti. Models with circular and eccentric orbits were considered. In all models, we started performing a maximum-likelihood fit to define the prior centers of the fit manually, then the MCMC model exploration was performed. Uniform bounds priors were utilized for: the planet’s period (between 0 and 20 days with initial guess at 6 days); the planet’s semi-amplitude (between 0 and 20 km s-1 with initial guess 6 m s-1); the SPIRou uncorrelated noise (between 0 and 10 m s-1 with initial guess 1 m s-1; and for (2459648.010 days and initial guess 2459648.0). The slope, dvdt, and the curvature, curv, were set equal to 0. We fed RadVel with the median-subtracted RV points, thus, the initial guess for the velocity zero-point was set to zero. The eccentricity was left to vary from 0 to 0.99. We used a stellar mass of 0.550.02 M⊙. The MCMC RadVel modeling used 50 walkers, 10000 steps, 8 ensembles and a minimum autocorrelation factor of 40.
Table 9 show the results of the RadVel AIC tests. In the km s-1 data the circular orbit is ”somewhat disfavored” with respect to the eccentric orbit. In the wPCA data the eccentric solution is favored, but the circular orbit is ”nearly indistinguishable”. In the Wapiti data the circular solution is favored, but the eccentric solutions is ”nearly indistinguishable”. The eccentricity distributions retrieved by RadVel are for the km s-1 data, for the wPCA data and for the Wapiti data. In all data-sets, the eccentricity is constrained to be lower than 0.4. As the circular and eccentric orbits provides statistically nearly indistinguishable solutions and the center of the eccentricity distributions is below 2 away from zero, we favor a circular orbit model. It has a lower number of free parameters and provides a good description of the data.
Table 10 summarizes the retrieved orbital and planet parameters for both the circular orbit model. Fig. 8 displays the phase-folded SPIRou RVs together with the best-fit circular orbit model for each of the SPIRou RV reductions, and the MCMC corner plots for the planet’s mass and semi-major axis. The RadVel models further confirm the detection of Gl 410b, as for all SPIRou data-sets the model without planet is “ruled-out”. Across all the SPIRou data-sets and models the period of the planet converges to a 6.02 days, thus a semi-mayor axis of the planet =0.0531 au. The wPCA and Wapiti data-sets give for the circular models similar results. The semi-amplitude retrieved in the LBL km s-1 data is larger, 5.30.7 m s-1, which translate to a of .
We note that the PCA methods filtering the LBL data are in fact decreasing the global dispersion of the all the RVs (not only those at km s-1) likely because of the diminution of the correlated noise and telluric correction residuals. However, the estimations of the planet mass with and without PCA correction methods are consistent within their 1 error bar. As the wPCA displays the second peak in the periodogram, and the time-series has the best correction of the km s-1 data, we favor the wPCA model. In the following, the wPCA circular orbital model is our baseline solution.
Taking the results altogether, SPIRou RVs analysis indicate that the candidate planet Gl 410b has a of Earth masses, orbits at a distance 0.05310.0006 astronomical units of its central star, with a period of 6.0200.004 days in a circular orbit. Gl 410b is a sub-Neptune planet.
6.2 Gl 410b properties retrieved from the SOPHIE RVs corrected by activity.
As discussed in Sect. 5.2, two methods were used to correct for the stellar activity RV jitter in the SOPHIE RV data and search for the RV signal of Gl 410b. In the first method, a RadVel Keplerian + GP model is fit to the SOPHIE RV data using the SOPHIE RV data alone. In the second method, a GP model trained using SPIRou’s longitudinal magnetic field measurements is fit to the SOPHIE RV time-series, and then a RadVel Keplerian model is fit to the RV residuals. In both methods, the Keplerian signal of Gl 410b is recovered at 6 days. The aim of this section is to check what are the planet properties derived, in particular , from the two activity correction methods and compare them to properties retrieved from SPIRou RVs.
When GPs are used to correct for the stellar activity RV jitter, there is uncertainty on how good the activity correction is, and therefore, how reliable the planet’s properties derived are. Using the planet properties derived from SPIRou wPCA RVs as benchmark, we aim to explore what activity correction method provides the closer match. As the SPIRou data suggests a circular orbit, for this analysis, we only modeled circular orbits. We limited the analysis to the 2021 to 2023 SOPHIE observations, as these data have the quasi-simultaneous SPIRou measurements.
| Model | RMS | BIC | AICc | AICs | |
| AICc Favored Model | |||||
| , , | 5 | 4.6 | 382 | 372 | 0.0 |
| Ruled Out | |||||
| , | 2 | 6.1 | 883 | 880 | 507 |
As previously discussed, for the RadVel Keplerian + GP model using the SOPHIE RV data alone, we used an uniform distribution priors for the planet’s orbital period (0 to 20 days, i.e. we leave RadVel free to find the period), the planet’s semi-amplitude (0.0 to 20.0 m s-1), the time of inferior conjunction tc (2459648.010 days) and the uncorrelated noise (0.0 to 15.0 m s-1).
We tested the same broad priors on the RadVel model of the RV data corrected by activity with the GP trained on . RadVel retrieved the expected the period of 6 days. However, to have a better determination of the planet parameters, we narrowed down the uniform priors to (5 to 7 days)111111 This is justified because in the SOPHIE RV-GP (trained on Bℓ) data-set, the periodogram shows a strong peak at 6.02 days and the 6 day period is retrieved using the broad prior.. tc (2459648.03 days) and (2.0 to 20.0 m s-1).
Table 4 presented already the results of the RadVel AIC tests for the Keplerian + GP model for the GP using the SOPHIE RV data alone. Table 7 describes the results of the AIC tests for the SOPHIE RV - GP trained on the time-series. For both activity correction methods, the RadVel AIC tests “rule-out” the model without a planet. Table 12 summarizes the orbital solutions found by RadVel for both activity methods. Both methods of activity filtering suggest the presence of a sub-Neptune mass planet with period 6.02 days and semimayor axis 0.0531 au.
| SOPHIEGP | SOPHIE-GP | |
| GP on RVSOPHIE alone | GP trained on | |
| Orbital parameters | ||
| [days] | ||
| [m s-1] | ||
| [JD] | ||
| Planet parameters | ||
| [M⊕] | ||
| [au] | ||
The SOPHIE RV RadVel GP + Keplerian model retrieves a of 7.6 m s-1 thus a of 14.5 M⊕. The SOPHIE - GP trained on time-series model retrieves a of m s-1 and of M⊕. While both methods provide a , thus a , consistent between them at the 1 level, the activity correction method that provides the closest match to the derived from the SPIRou RVs, is the correction by the GP trained on SPIRou’s . We display the model in Figure 9.
Our analysis suggests that the measurements of performed in the near-IR with SPIRou could be indeed a good activity proxy to correct for the stellar RV activity jitter in quasi-simultaneous measurements taken in the optical. The independent detection of the same signal in both SPIRou and SOPHIE data-sets gives solid support to a planet as being the responsible of the periodic variation seen in near-IR and optical RVs.
6.3 A joint model for Gl 410b using the SPIRou wPCA and the activity corrected SOPHIE data-sets
As the planet signal is recovered independently in the near-IR and in the optical, a joint model using both RV data-sets can be performed. To merge RV data from different wavelengths and instruments has risks and can be problematic, as instruments have different RV systematic noise and stellar activity is chromatic. Nevertheless, to check, we used RadVel and calculated a joint solution for the orbital parameters of Gl 410b using the wPCA SPIRou RVs and the GP activity corrected (using as activity proxy) SOPHIE RVs. We used an uniform distribution priors for the planet’s orbital period (0 to 20 days), the planet’s semi-amplitude (0.0 to 20.0 m s-1), the time of inferior conjunction tc (2459648.010 days) and the uncorrelated noise (0.0 to 10.0 m s-1) and (0.0 to 10.0 m s-1). We tested models with circular and eccentric orbits.
Table 9 summarises the results of the RadVel AIC tests. The model without planet is “ruled-out” by RadVel. The model with a planet in circular orbit is favoured, although a model with an eccentric orbit is “nearly indistinguishable”. We favor the circular solution because it describes correctly the data with fewer free parameters. We display the retrieved planet properties in Fig. 27 in the Appendix. The combination of the wPCA SPIRou and the activity corrected SOPHIE-GP data gives essentially the same planet properties as the wPCA SPIRou data alone. The only (small) differences are that a decrease in the period uncertainty from 0.004 to 0.003 days and a increase from M⊕ to M⊕ ( m s-1 to m s-1) in the combined data. The larger dispersion of the SOPHIEGP RVs with respect of wPCA SPIRou RVs could be responsible of the small increase on the planet’s semi-amplitude. The uncertainty in the period becomes smaller in the combined data-set because in both data-sets the planet signal is time and phase coherent.
| Model | RMS | BIC | AICc | AICs | |
| AICc Favored Model | |||||
| , , | 7 | 5.6 | 1403 | 1380.3 | 0.0 |
| Nearly Indistinguinshable | |||||
| , , , | 5 | 5.6 | 1410 | 1380.6 | 0.34 |
| Ruled Out | |||||
| , | 2 | 6.6 | 3308 | 3295 | 1915 |
7 Search for additional signals within the SPIRou wPCA data.
As presented in Sect. 4.2 the GLS periodogram of the wPCA SPIRou RVs shows in addition to the 6.02 day peak a second peak at 18.76 days with FAPboostrap=1.6%. The presence of additional planetary signals in the data is encouraging as it is known from transit searches that M-dwarfs often exhibit multi-planetary systems. In this section we investigate the possibility of having additional planetary signals within the wPCA SPIRou RVs.
7.1 A planet candidate at P=18.7 days ?
If the 6.02 day Keplerian signal of Gl 410b (Table 10) is subtracted from the wPCA SPIRou time-series and a new GLS periodogram is calculated on the residuals, we find that the signal at 18.76 days is still present. Furthermore, its FAPboostrap level improves to the 0.3% level (Fig. 10 middle panel). This provides tentative evidence in favour for the presence of a second planet.
To further test the plausibility of the existence of the signal at 18.7 days, we calculated a RadVel MCMC run using a two-planets model. For this test we set broad uniform priors for (0 to 20 days), (0 to 30 days), and and (0 and 20 m s-1). Circular and eccentric orbits models were calculated. The model priors and initial guesses are summarized in Table 17 in the Appendix. Our goal was to determine which planet solution is favored by RadVel and compare the periods blindly retrieved by RadVel with those of the periodogram analysis.
The RadVel AIC model comparaison is presented in Table 15 in the Appendix. RadVel finds that is more likely the two-planet solution over the one-planet solution. In fact, the models with one planet are “ruled-out” by RadVel. RadVel favors a model which planet b has an eccentric orbit and planet c a circular orbit; and a model with circular orbits for both b and c. However, both models are “nearly indistinguishable”. We favor the model in which the two-planets are in circular orbits.
In the Appendix, we provide the corner plots of the 2-planet in circular orbits model (Fig 28), a plot showing the phase-folded RVs and the best fitting 2-planet circular solution (Fig 29), and a table summarising the orbital parameters and physical properties retrieved (Table 18).
RadVel indicates for the candidate signal at 18.7 days a = M⊕ ( m s-1 and a period =18.760.03 days (0.1130.001 au). This period, retrieved blindly by RadVel, corresponds to the peak in the periodogram once the RV signal of Gl 410b is removed from the data (middle panel of Fig. 10). The properties of Gl 410b are similar to those retrieved in the one-planet model, slightly decreases to 7.91.2 M⊕, which is within the 1 error of the one-planet model value. The bottom panel of Fig. 10, displays the periodogram of the residuals of SPIRou wPCA RVs after subtraction of the signal of Gl 410b and the candidate signal at 18.7 days. No peaks with significance higher than FAP=10% are left in the periodogram. We tested a RadVel model on the combined SPIRou wPCA and SOPHIEGP data-set (not shown), in this data also the two-planets in circular orbits model was favored by the AIC tests.
Gl 410 has thus tantalising evidence for hosting two sub-Neptune planets in circular orbits close to 3:1 resonance. For curiosity, using the RadVel posterior distributions of the eccentric orbit models, we explored the stability of the system (Sect. G in the Appendix). We found that the system is safely stable for circular orbits and the nominal eccentricity values.
7.2 -periodogram / apodized sines analysis
The results of the GLS periodogram and the 2-planet Radvel modeling are promising. To perform an independent assessment on the existence of periodicities within the SPIRou wPCA data-set, we performed a and apodize sine periodogram (ASP) analysis. The periodogram searches for several signals simultaneously, and the ASP tests the consistency in frequency, phase, and amplitude of the candidate signals. We used the periodogram technique, as defined in Hara et al. (2017). This tool is based on a sparse recovery technique called the basis pursuit algorithm (Chen et al., 1998). It aims to find a representation of the RV time series as a sum of a small number of sinusoids whose frequencies are in the input grid. It outputs a figure which has a similar aspect as a regular periodogram, but with fewer peaks due to aliasing. The peaks can be assigned a FAP, whose interpretation is close to the FAP of a regular periodogram peak.
The periodogram takes several parameters in input, in particular a list of vectors, or predictor, which are fitted linearly along with the search for periodic signals. We chose for this linear base model several stellar activity indicators available within the SPIRou data: the longitudinal magnetic field () measurement, the LBL projection onto the 2nd derivative of the spectrum (d2v), the LBL projection onto the 3rd derivative of the spectrum (d3v), the LBL differential line width (dLW), and LBL the differential temperature measurement (dTEMP) with respect to a K template. The LBL activity indicators are calculated at the same time that the RVs (Artigau et al., 2022, 2024). We also added a periodic signal at the sidereal day, to account for apparent systematics at this period. The periodogram also necessitates to assume a certain form of the noise model. Following the method of Hara et al. (2020), we tryed a grid of noise models, and rank them with cross-validation and Bayesian evidence, computed with the Laplace approximation. The detailed analysis is presented in Appendix H.1.
In Figure 11, we represent the periodogram of the noise model with the highest cross-validation score. We find clear evidence for the 6.02 days signal, and a FAP of . The 18.7 day signal is also present with a FAP. We note the presence of an additional 2.99 day signal with a FAP . Finally, we find a marginally significant signal at 1.3 days. The signals at 6, 18 and 3 days are consistently found by the periodograms computed: they present peaks in 100% of the top 20% of the models with highest cross validation scores.
To further refine the analysis, we tested whether the candidate signals are strictly periodic, or if their frequency, phase and amplitudes show signs of variation. For this, we employed the method of Hara et al. (2022), which consists in fitting apodized sinusoidal signals to the dataset. These are wavelet-like signals, consisting of a sine multiplied by a Gaussian function, localized in time. In Appendix H.1, we show that there is good evidence that the three signals found by the periodogram are strictly periodic.
In summary, besides the clear 6.02 day planet, we find in the and ASP analysis statistically significant candidates at 18.7 and 2.99 d, which are compatible with purely periodic signals. Let us note that the proximity of the 2.99 d planet to a 2:1 mean motion resonance with the 6 day planet further supports that indeed is a planet. Nonetheless, we do not label the two signals as confirmed planets, as their statistical significance is not completely sufficient.
To conclude this section, we would like to insist that the detections of the signals at 18.7 and 2.99 days, although they pass several statistical tests, at the present are candidate signals because they are retrieved only on the wPCA SPIRou data. Further measurements and improvements in the methods to recover the data in the —Vtot—¡10 km s-1 region are required to fully exploit all the measurements taken and ultimately confirm or refute the planetary nature of these signals.
7.3 Planet detection limits at long periods from the SPIRou wPCA radial velocities
We computed planet detection limits from the SPIRou RVs performing an injection recovery test. The method determines the upper limits by injecting signals in the time series at given periods, calculating the GLS periodograms and determining the maximum amplitude that can still be missed with a given probability. For each period, the upper limit is given by the highest amplitude (at the worst phase) that could be present in the time series without creating a peak above the 1% FAP level in the periodogram. The method consists in injecting for each period a sinusoidal Keplerian signal (i.e. circular orbit) at 12 equi-spaced phases. The amplitude (i.e., the planet signal strength) is then increased until the corresponding peak in the GLS periodogram reaches the 1% FAP level. The 1% FAP confidence level was computed from the levels of the highest peaks of 1000 permuted data sets, and the 1% upper limits’ function was obtained by sorting the power of peaks of all periodograms at each period. We tested periods from 0.8 to 10000 days.
Figure 12 displays the projected mass as a function of the period at which 99% and 75% of the planets injected would be detected in the GLS periodogram of the SPIRou data. The degraded upper limits for periods around 1, 2, 180, 360 days are due to the data sampling and an unavoidable consequence of observing at night from a single location and that SPIRou is only mounted in bright time. Our SPIRou RV data would have recovered with a 99% confidence planets with masses larger than 0.1, 1.0 and 4.7 MJ at a periods of 100, 2560 and 5000 days (0.35, 3 and 5 au) respectively.
Kervella et al. (2022) perform a study aimed at detecting companions of Hipparcos catalog stars based on the proper motion anomaly (PMa) derived from Hipparcos and Gaia measurements. Gl 410 is part of that study, it displays a PMa S/N of 0.13. Kervella et al. (2022) derive companion mass of 0.02 MJ (6.3 M⊕) in a 3 to 5 au orbit. The mass of this companion candidate is significantly below the sensitivity capability of our current SPIRou RV measurements.
8 TESS photometry analysis
The TESS photometry time-series is shown in Fig. 13. One can clearly observe the modulation of the TESS light-curve by the stellar rotation. However, it is not possible to measure the rotation period with high accuracy, because the time span of the observations (27 days) does not fully cover two complete rotation periods. No transits were identified by the TESS automatic pipeline. To further search for transit events, we applied a procedure customized for Gl 410. After removing the outliers using a 3 clipping procedure, we used the package Wotan (Hippke et al., 2019) to flatten the light curve and to remove the effects of variability due to stellar rotation. The second panel of Fig. 13 displays the resultant detrended light curve, it has a scatter of 403 ppm. Using the transit least squares algorithm of Hippke & Heller (2019), we searched for periodic transit events in the detrended light-curve, in particular at the orbital period of planet b. The algorithm retrieves periodic signals at 3.5 and 7 days with a signal detection efficiency (SDE) of the order of 5, which is not high enough to claim a detection (Fig. 14). Further inspection of those signals reveal they are artefacts. The signals are close to the harmonics of the stellar rotation period and likely are related to residuals of the detrending procedure. No signal is detected at period of 6.02 days, suggesting that Gl 410b is not transiting.
8.1 Transit detection limits
To determine the detection limits in the TESS data, we conducted an injection recovery test of transit events. Following Cortés-Zuleta et al. (2025), using the stellar parameters of Gl 410 and the semi-mayor axis of 0.053 au, we injected 50 000 transits generated with the software batman (Kreidberg, 2015), on which the orbital inclination and planet radius are drawn randomly from uniform distributions: between 85∘ and 90∘ for the inclination and from 0.1 to 10 for the radius. For each synthetic transit, we computed the box least squares (BLS) periodogram131313https://docs.astropy.org/en/stable/timeseries/bls.html implemented in astropy and measured the power of the signal at a period of 6.02 days. Then, we measured the fraction of transit events that are detectable with the BLS periodogram (i.e., power 1000). The results of the injection test are shown in Fig. 15. We find that planets with radius smaller than 1.8 R⊕ are not detectable in the current TESS data, as well as planets in orbits with inclinations lower than 87 deg.
9 Discussion
9.1 Gl 410b in context
Gl 410b sums up to the growing list of sub-Neptune mass planets detected around low-mass stars ( M⊙). In order to situate the Gl 410 planetary system in context, we will concentrate here in discussing the planet detection statistics of the stars in the 0.5 - 0.6 M⊙ stellar mass bin, the stars with a mass equal to that of Gl 410. This stellar mass range has at the time of writing (01.07.2024) according to the NASA exoplanets catalogue141414https://exoplanetarchive.ipac.caltech.edu 212 planet candidates detected (of 5678 exoplanets). Several of those planets are awaiting further characterization, but already 140 have a reported radius measurement, 83 have a planetary mass measurement and 30 have a mass constraint (i.e., ).
The period is always constrained both in transit and RV planet detections. However, in both methods, detection statistics are strongly positively biased to short periods, as planets in short periods are the easiest to detect. At the time of writing, in the M⊙ stellar mass bin, from 212 planet candidates, there are 161 detections at 50 days, with 80 % of those detections at 15 days. These numbers illustrate the observational biases mentioned. To perform a detailed study correcting for biases, it is out of the scope of this discovery paper, however, we can point out, that planets in a 6 days period, such as Gl 410b, are not unusual in the M⊙ stellar mass bin.
Concerning the planet mass, we are interested in discussing how abundant or scarce are planets in the mass range of Gl 410b (8.41.3 M⊕) in orbits with periods 50 days around stars in the M⊙ stellar mass bin. We show in Fig. 16 the histograms of the planet mass distribution for a days (top panel) and for between 10 and 50 days (bottom panel). The histogram includes both mass measurements and lower-limits. In the sub-sample with days, the most abundant planets are those with masses 10 M⊕ with 40% of the detections, followed by planets with mass between 1/3 and 1 MJ with 23% of the detections. Detection biases make planets with masses 10 M⊕ the hardest to detect; the large fraction of detections of planets in the low mass range, means that they should be intrinsically the most numerous. The histograms show a significant decrease on the frequency of planets for the mass bin between 17 M⊕ (Neptune mass) and 100 M⊕ ( 1/3 MJ). As planets in this mass range are more easily detected by RVs than planets with 10 M⊕, the decrease is real and it is not a detection bias. We find that Gl 410b is located in the bin with the highest probability of detection.
In the context of the sub-sample of planets with periods between 10 and 50 days, we find that: i) the highest fraction of detections is the mass bin between 5 and 10 M⊕; ii) there is a decrease in the number of detections of planets with 5 M⊕; and iii) there are very few detections of giant planets. The smaller fraction of low-mass planets could be an observational bias given the fact that to detect planets with 5 M⊕ around M dwarfs at longer periods it is likely harder because the amplitude becomes smaller and it could be comparable to the activity jitter. The very small fraction of giant planets, in contrast, is not an observational bias, must be real, as giant planets are easier to detect by radial velocities. The tentative planet candidate signal at 18.7 days would thus belong to the category of the most frequent planets detected.
The very fact that Gl 410b is located in the bins with the highest number of detections, reinforces the current view that sub-Neptune mass planets are abundant around M dwarfs. The decrease of the occurence rate with increasing planetary mass, it is a property already highlighted in studies covering the entire M dwarf mass domain e.g., Bonfils et al. (2013), Sabotta et al. (2021), Mignon et al. (2025). Current statistics raise the question of why planets with masses between Neptune and 1/3 MJ are scarce at periods shorter than 50 days around stars whose mass is in the M⊙ mass bin. Whether this could be a consequence of the planet formation process, or a product of the interaction of planetary atmospheres with energetic radiation and stellar wind particles due to the high activity levels of M dwarfs is an open question.
As Gl 410b is not transiting, we don’t have a stringent constraint in their radius. From the sample of planets in M⊙ stellar mass bin, days and masses similar as Gl 410b (7 to 10 M⊙) there is at present only a few planets with measurements of the radius for example TOI-1725b ( R⊕, =9.9 M⊕, Essack et al., 2023), and TOI-2018b ( R⊕, =9.2 M⊕, Dai et al., 2023).
9.2 Stellar activity and the evolution of Gl 410b
Let us conclude this section briefly discussing the implications of stellar activity in the atmospheric evolution of Gl 410b151515 Pending the confirmation of the planet at 2.99 days with new measurements, we not mention it in this discussion. However, the arguments exposed here for Gl 410b would be applicable to this planet given its shorter period.. As presented in Section 2, Gl 410 is an active star. Activity is the manifestation of the presence of stellar magnetic fields. Magnetic fields heat the chromosphere and coronae and drive stellar winds (e.g. Cranmer & Winebarger, 2019). High-energy radiation from stellar flares and X-ray/EUV coronal emissions, and particle fluxes from winds and coronal mass ejections interact with the magnetospheres and the atmospheres of planets in orbits close to the star driving thermal and non-thermal atmospheric escape processes (for recent reviews see Vidotto, 2022; Kubyshkina, 2024).
Gl 410b is located very close to their host star at 0.05 au. It is therefore expected to have an active interaction with the high-energy radiation and the stellar wind particle fluxes from Gl 410. Using the radius and given in Table 1, we obtain that Gl 410b receive 20.4 Earth’s insulation. It is out of the scope of this discovery paper to perform a detailed study on the conditions on which a planet with 9 Earth masses could retain its atmospheres in case of close proximity with an active M-dwarf such as Gl 410, but it is interesting topic that deserves further study. Furthermore, it would be of extreme interest to investigate the influence that stellar irradiation and activity could have in the chemical composition of the planet’s atmosphere and the potential development of abundance anomalies due to the escape of volatile species (e.g., Louca et al., 2023) in single and multi-planetary systems (e.g., Acuña et al., 2022).
10 Conclusion
In this paper, we presented a radial-velocity study of Gl 410, a nearby (=12 pc) early M dwarf with a mass of 0.55 M⊙ and an age of 480150 Myr 161616This work, see Appendix A.. We have monitored the star in the near-IR with the spectropolarimeter and velocimeter SPIRou and in the optical with the velocimeter SOPHIE. The SPIRou RVs show a robust periodic signal at =6.0200.004 days. The signal is recovered in the raw LBL data and in the two PCA methods used to correct the LBL RVs for systematics and telluric correction residuals. In our baseline solution, the circular orbit model to the SPIRou wPCA data-set (Table 10), the signal has a semi-amplitude of 4.40.7 m s-1, which indicates a sub-Neptune planet with sin of 8.41.3 Earth masses at a 0.05310.0006 au. The same Keplerian signal is recovered in quasi-simultaneous SOPHIE RV measurements after correction for stellar activity using a GP that has been trained on the measurements of the longitudinal magnetic field () obtained with SPIRou. The independent detection of Gl 410b in the SOPHIE RVs provide further evidence for the Keplerian nature of the periodic signal.
We investigated the TESS archive and searched whether Gl 410b could be a transiting planet. We find no evidence for the transit of Gl 410b. A recovery analysis on the TESS photometry indicates that the transit would have been detected if Gl 410b’s inclination would be higher than 87 deg and its radius larger than 1.8 R⊕. Gl 410b is located very close to their host star, it receives 20.4 Earth’s insulation. It would be of great interest to investigate the influence that stellar irradiation and activity could have in the chemical composition of the planet’s atmosphere and in potential atmospheric escape processes.
We find within the SPIRou wPCA RVs tentative evidence for two additional planetary signals at 2.99 and 18.7 days. The 18.7 day signal (consistent with a sub-Neptune planet) is directly seen in the GLS periodograms and it is recovered in a Radvel two-planet model with broad priors for the period. The 2.99 day signal emerges together with the 18.7 day signal in an and apodize sine periodogram (ASP) analysis which takes into account stellar activity indicators measured with SPIRou and a model for the correlated noise. The planet at 2.99 days could be in 2:1 mean motion resonance with the 6.02 days planet.
The detection of Gl 410b shows that infrared RVs are a powerful tool to detect planets around active stars, objects that are often excluded from searches in optical RVs. Care should be taken however to correct/filter systematics generated by residuals of the telluric correction or small structures in the detector plane. The LBL technique combined with PCA offers a promising way to reach this objective. They enabled us to use the RV measurements acquired when the star is at km s-1 of the center of telluric absorption lines, which is important to detect additional planetary signals. Further monitoring of Gl 410 is necessary.
Acknowledgements.
Based on observations obtained at the Canada-France-Hawaii Telescope (CFHT) which is operated from the summit of Maunakea by the National Research Council of Canada, the Institut National des Sciences de l’Univers of the Centre National de la Recherche Scientifique of France, and the University of Hawaii. Based on observations obtained with SPIRou, an international project led by Institut de Recherche en Astrophysique et Planétologie, Toulouse, France. Based on observations obtained with the spectrograph SOPHIE at the Observatoire de Haute-Provence (OHP) in France, operated by Institut National des Sciences de l’Univers of the Centre National de la Recherche Scientifique of France. We acknowledge funding from the French ANR under contract number ANR18CE310019 (SPlaSH). This work is supported by the French National Research Agency in the framework of the Investissements d’Avenir program (ANR-15-IDEX-02), through the funding of the “Origin of Life” project of the Grenoble-Alpes University. JFD acknowledges funding from the European Research Council (ERC) under the H2020 research & innovation program (grant agreement #740651 NewWorlds). This publication makes use of the Data & Analysis Center for Exo-planets (DACE), which is a facility based at the University of Geneva (CH) dedicated to extrasolar planets data visualization, exchange and analysis. DACE is a platform of the Swiss National Centre of Competence in Research (NCCR) PlanetS, federating the Swiss expertise in Exoplanet research. The DACE platform is available at https://dace.unige.ch. The project leading to this publication has received funding from the french government under the ”France 2030” investment plan managed by the French National Research Agency (reference : ANR-16-CONV-000X / ANR-17-EURE-00XX) and from Excellence Initiative of Aix-Marseille University - A*MIDEX (reference AMX-21-IET-018). This work was supported by the ”Programme National de Planétologie” (PNP) of CNRS/INSU. E.M. acknowledges funding from FAPEMIG under project number APQ-02493-22 and research productivity grant number 309829/2022-4 awarded by the CNPq, Brazil. This work has made use of data from the European Space Agency (ESA) mission Gaia (https://www.cosmos.esa.int/gaia), processed by the Gaia Data Processing and Analysis Consortium (DPAC, https://www.cosmos.esa.int/web/gaia/dpac/consortium). Funding for the DPAC has been provided by national institutions, in particular the institutions participating in the Gaia Multilateral Agreement. The observations at the Canada-France-Hawaii Telescope were performed with care and respect from the summit of Maunakea which is a significant cultural and historic site.References
- Acuña et al. (2022) Acuña, L., Lopez, T. A., Morel, T., et al. 2022, A&A, 660, A102
- Ammler-von Eiff & Guenther (2009) Ammler-von Eiff, M. & Guenther, E. W. 2009, A&A, 508, 677
- Angus et al. (2019) Angus, R., Morton, T., & Foreman-Mackey, D. 2019, The Journal of Open Source Software, 4, 1469
- Artigau et al. (2014) Artigau, É., Astudillo-Defru, N., Delfosse, X., et al. 2014, in Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 9149, Observatory Operations: Strategies, Processes, and Systems V, ed. A. B. Peck, C. R. Benn, & R. L. Seaman, 914905
- Artigau et al. (2024) Artigau, É., Cadieux, C., Cook, N. J., et al. 2024, AJ, 168, 252
- Artigau et al. (2022) Artigau, É., Cadieux, C., Cook, N. J., et al. 2022, AJ, 164, 84
- Astudillo-Defru et al. (2015) Astudillo-Defru, N., Bonfils, X., Delfosse, X., et al. 2015, A&A, 575, A119
- Bailer-Jones et al. (2021) Bailer-Jones, C. A. L., Rybizki, J., Fouesneau, M., Demleitner, M., & Andrae, R. 2021, AJ, 161, 147
- Bellotti (2023) Bellotti, S. 2023, PhD thesis, IRAP
- Bellotti et al. (2024) Bellotti, S., Morin, J., Lehmann, L. T., et al. 2024, A&A, 686, A66
- Bertaux et al. (2014) Bertaux, J. L., Lallement, R., Ferron, S., Boonne, C., & Bodichon, R. 2014, A&A, 564, A46
- Boisse et al. (2009) Boisse, I., Moutou, C., Vidal-Madjar, A., et al. 2009, A&A, 495, 959
- Bonfils et al. (2013) Bonfils, X., Delfosse, X., Udry, S., et al. 2013, A&A, 549, A109
- Bouchy et al. (2013) Bouchy, F., Díaz, R. F., Hébrard, G., et al. 2013, A&A, 549, A49
- Bouchy et al. (2017) Bouchy, F., Doyon, R., Artigau, É., et al. 2017, The Messenger, 169, 21
- Bouchy et al. (2009) Bouchy, F., Hébrard, G., Udry, S., et al. 2009, A&A, 505, 853
- Bouma et al. (2023) Bouma, L. G., Palumbo, E. K., & Hillenbrand, L. A. 2023, ApJ, 947, L3
- Cadieux et al. (2022) Cadieux, C., Doyon, R., Plotnykov, M., et al. 2022, AJ, 164, 96
- Carmona et al. (2023) Carmona, A., Delfosse, X., Bellotti, S., et al. 2023, A&A, 674, A110
- Chauvin (2024) Chauvin, G. 2024, Comptes Rendus Physique, 24, 139
- Chen et al. (1998) Chen, S. S., Donoho, D. L., & Saunders, M. A. 1998, SIAM JOURNAL ON SCIENTIFIC COMPUTING, 20, 33
- Chen et al. (2020) Chen, X., Ge, Z., Chen, Y., et al. 2020, The Astrophysical Journal, 889, 157
- Cook et al. (2022) Cook, N. J., Artigau, É., Doyon, R., et al. 2022, arXiv e-prints, arXiv:2211.01358
- Cortés-Zuleta et al. (2025) Cortés-Zuleta, P., Boisse, I., Ould-Elhkim, M., et al. 2025, A&A, 693, A164
- Cranmer & Winebarger (2019) Cranmer, S. R. & Winebarger, A. R. 2019, ARA&A, 57, 157
- Cretignier et al. (2023) Cretignier, M., Dumusque, X., Aigrain, S., & Pepe, F. 2023, A&A, 678, A2
- Cristofari et al. (2022) Cristofari, P. I., Donati, J. F., Masseron, T., et al. 2022, MNRAS, 516, 3802
- Cristofari et al. (2023) Cristofari, P. I., Donati, J. F., Moutou, C., et al. 2023, MNRAS, 526, 5648
- Curtis et al. (2020) Curtis, J. L., Agüeros, M. A., Matt, S. P., et al. 2020, ApJ, 904, 140
- Cutri et al. (2003) Cutri, R. M., Skrutskie, M. F., van Dyk, S., et al. 2003, VizieR Online Data Catalog: 2MASS All-Sky Catalog of Point Sources (Cutri+ 2003), VizieR On-line Data Catalog: II/246. Originally published in: University of Massachusetts and Infrared Processing and Analysis Center, (IPAC/California Institute of Technology) (2003)
- Czesla et al. (2019) Czesla, S., Schröter, S., Schneider, C. P., et al. 2019, PyA: Python astronomy-related packages
- Dai et al. (2023) Dai, F., Schlaufman, K. C., Reggiani, H., et al. 2023, AJ, 166, 49
- Donati et al. (2020) Donati, J. F., Kouach, D., Moutou, C., et al. 2020, MNRAS, 498, 5684
- Donati et al. (2023) Donati, J. F., Lehmann, L. T., Cristofari, P. I., et al. 2023, MNRAS, 525, 2015
- Donati et al. (2008) Donati, J. F., Morin, J., Petit, P., et al. 2008, MNRAS, 390, 545
- Donati et al. (1997) Donati, J. F., Semel, M., Carter, B. D., Rees, D. E., & Collier Cameron, A. 1997, MNRAS, 291, 658
- Dressing & Charbonneau (2013) Dressing, C. D. & Charbonneau, D. 2013, ApJ, 767, 95
- Dressing & Charbonneau (2015) Dressing, C. D. & Charbonneau, D. 2015, ApJ, 807, 45
- Essack et al. (2023) Essack, Z., Shporer, A., Burt, J. A., et al. 2023, AJ, 165, 47
- Fekel & Henry (2000) Fekel, F. C. & Henry, G. W. 2000, AJ, 120, 3265
- Foreman-Mackey (2015) Foreman-Mackey, D. 2015, George: Gaussian Process regression, Astrophysics Source Code Library, record ascl:1511.015
- Foreman-Mackey et al. (2013) Foreman-Mackey, D., Conley, A., Meierjurgen Farr, W., et al. 2013, emcee: The MCMC Hammer, Astrophysics Source Code Library, record ascl:1303.002
- Fouqué et al. (2023) Fouqué, P., Martioli, E., Donati, J. F., et al. 2023, A&A, 672, A52
- Fulton et al. (2018) Fulton, B. J., Petigura, E. A., Blunt, S., & Sinukoff, E. 2018, PASP, 130, 044504
- Gagné et al. (2021) Gagné, J., Faherty, J. K., Moranta, L., & Popinchalk, M. 2021, ApJ, 915, L29
- Gagné et al. (2020) Gagné, J., Faherty, J. K., & Popinchalk, M. 2020, Research Notes of the American Astronomical Society, 4, 92
- Gaia Collaboration et al. (2023) Gaia Collaboration, Vallenari, A., Brown, A. G. A., et al. 2023, A&A, 674, A1
- Gaidos et al. (2023) Gaidos, E., Claytor, Z., Dungee, R., Ali, A., & Feiden, G. A. 2023, MNRAS, 520, 5283
- Gladman (1993) Gladman, B. 1993, Icarus, 106, 247
- Hara et al. (2020) Hara, N. C., Bouchy, F., Stalport, M., et al. 2020, A&A, 636, L6
- Hara et al. (2017) Hara, N. C., Boué, G., Laskar, J., & Correia, A. C. M. 2017, MNRAS, 464, 1220
- Hara et al. (2019) Hara, N. C., Boué, G., Laskar, J., Delisle, J. B., & Unger, N. 2019, MNRAS, 489, 738
- Hara et al. (2022) Hara, N. C., Delisle, J.-B., Unger, N., & Dumusque, X. 2022, A&A, 658, A177
- Hébrard et al. (2016) Hébrard, É. M., Donati, J. F., Delfosse, X., et al. 2016, MNRAS, 461, 1465
- Heidari (2022) Heidari, N. 2022, Theses, Université Côte d’Azur ; Shahid Beheshti University (Tehran)
- Hippke et al. (2019) Hippke, M., David, T. J., Mulders, G. D., & Heller, R. 2019, AJ, 158, 143
- Hippke & Heller (2019) Hippke, M. & Heller, R. 2019, A&A, 623, A39
- Hobson et al. (2018) Hobson, M. J., Díaz, R. F., Delfosse, X., et al. 2018, A&A, 618, A103
- Hsu et al. (2020) Hsu, D. C., Ford, E. B., & Terrien, R. 2020, MNRAS, 498, 2249
- Jenkins et al. (2016) Jenkins, J. M., Twicken, J. D., McCauliff, S., et al. 2016, in Proc. SPIE, Vol. 9913, Software and Cyberinfrastructure for Astronomy IV, 99133E
- Kasper et al. (2021) Kasper, M., Cerpa Urra, N., Pathak, P., et al. 2021, The Messenger, 182, 38
- Kervella et al. (2022) Kervella, P., Arenou, F., & Thévenin, F. 2022, A&A, 657, A7
- King et al. (2003) King, J. R., Villarreal, A. R., Soderblom, D. R., Gulliver, A. F., & Adelman, S. J. 2003, AJ, 125, 1980
- Koen et al. (2010) Koen, C., Kilkenny, D., van Wyk, F., & Marang, F. 2010, MNRAS, 403, 1949
- Kreidberg (2015) Kreidberg, L. 2015, PASP, 127, 1161
- Kubyshkina (2024) Kubyshkina, D. 2024, Planetary atmospheres Through Time: Effects of Mass Loss and Thermal Evolution
- Louca et al. (2023) Louca, A. J., Miguel, Y., Tsai, S.-M., et al. 2023, MNRAS, 521, 3333
- Mahadevan et al. (2014) Mahadevan, S., Ramsey, L. W., Terrien, R., et al. 2014, in Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 9147, Ground-based and Airborne Instrumentation for Astronomy V, ed. S. K. Ramsay, I. S. McLean, & H. Takami, 91471G
- Ment & Charbonneau (2023) Ment, K. & Charbonneau, D. 2023, AJ, 165, 265
- Mignon et al. (2025) Mignon, L., Delfosse, X., Meunier, N., et al. 2025, arXiv e-prints, arXiv:2502.06553
- Montes et al. (2001) Montes, D., López-Santiago, J., Gálvez, M. C., et al. 2001, MNRAS, 328, 45
- Mulders et al. (2015) Mulders, G. D., Pascucci, I., & Apai, D. 2015, ApJ, 814, 130
- Nelson et al. (2020) Nelson, B. E., Ford, E. B., Buchner, J., et al. 2020, AJ, 159, 73
- Ould-Elhkim et al. (2023) Ould-Elhkim, M., Moutou, C., Donati, J. F., et al. 2023, A&A, 675, A187
- Perruchot et al. (2008) Perruchot, S., Kohler, D., Bouchy, F., et al. 2008, in Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 7014, Ground-based and Airborne Instrumentation for Astronomy II, ed. I. S. McLean & M. M. Casali, 70140J
- Petit et al. (2018) Petit, A. C., Laskar, J., & Boué, G. 2018, A&A, 617, A93
- Pinamonti et al. (2022) Pinamonti, M., Sozzetti, A., Maldonado, J., et al. 2022, A&A, 664, A65
- Quirrenbach et al. (2018) Quirrenbach, A., Amado, P. J., Ribas, I., et al. 2018, in Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 10702, Ground-based and Airborne Instrumentation for Astronomy VII, ed. C. J. Evans, L. Simard, & H. Takami, 107020W
- Reiners et al. (2022) Reiners, A., Shulyak, D., Käpylä, P. J., et al. 2022, A&A, 662, A41
- Ricker et al. (2014) Ricker, G. R., Winn, J. N., Vanderspek, R., et al. 2014, in Society of Photo-Optical Instrumentation Engineers (SPIE) Conference Series, Vol. 9143, Space Telescopes and Instrumentation 2014: Optical, Infrared, and Millimeter Wave, ed. J. Oschmann, Jacobus M., M. Clampin, G. G. Fazio, & H. A. MacEwen, 914320
- Sabotta et al. (2021) Sabotta, S., Schlecker, M., Chaturvedi, P., et al. 2021, A&A, 653, A114
- Schöfer et al. (2019) Schöfer, P., Jeffers, S. V., Reiners, A., et al. 2019, A&A, 623, A44
- Silva Aguirre et al. (2018) Silva Aguirre, V., Bojsen-Hansen, M., Slumstrup, D., et al. 2018, MNRAS, 475, 5487
- Smith et al. (2012) Smith, J. C., Stumpe, M. C., Van Cleve, J. E., et al. 2012, PASP, 124, 1000
- Stumpe et al. (2014) Stumpe, M. C., Smith, J. C., Catanzarite, J. H., et al. 2014, PASP, 126, 100
- Stumpe et al. (2012) Stumpe, M. C., Smith, J. C., Van Cleve, J. E., et al. 2012, PASP, 124, 985
- Vidotto (2022) Vidotto, A. A. 2022, arXiv e-prints, arXiv:2211.15396
- Zechmeister & Kürster (2009) Zechmeister, M. & Kürster, M. 2009, A&A, 496, 577
Appendix A On the age of Gl 410
Age determination using gyrochronology rely on empirical relations between the rotation period and the effective temperature or photometry of stars in clusters of known ages. The age of 0.890.1 Gyr for Gl 410 derived in Fouqué et al. (2023) was determined using the methods described in Gaidos et al. (2023) which are calibrated on the empirical relations of Curtis et al. (2020). In Fig. 17, we display the vs diagram from Curtis et al. (2020), indicating the location of Gl 410 in solid lines. At the Gl 410, the age-rotation relations of Praesepe (670 Myr) and NGC 6811 (1 Gyr) sequences overlap. The rotational period of Gl 410 locates the star below the Praesepe sequence, which suggests that Gl 410 could be younger than 670 Myr. The Gaidos et al. (2023) work targeted the vast majority of planet hosts which are middle-aged and did not use calibrations below the Praesepe age so by definition no claim for a younger age could be done.
To investigate the possibility of a younger age for Gl 410, we estimated its age using two alternative publicly available gyrochronology codes: stardate (Angus et al. 2019) and gyro-inter (Bouma et al. 2023). Both codes converge in a younger age estimate for Gl 410: stardate suggests an age of 4932 Myr; gyro-inter indicates an age of 326 Myr. Furthermore, we checked the position of Gl 410 in the color-magnitude diagram (CMD). The right panel of Fig. 17 displays the Gaia DR3 (Gaia Collaboration et al. 2023) CMD for Gl 410, field stars, together with empirical CMD sequences from various young associations (Gagné et al. 2021). Gl 410 is located right in between the 400 and 580 Myr sequences. Interpolating in log space, an age of about 480 Myr is obtained for Gl 410. Note however, that the scatter around these empirical sequences is non negligible, and that at younger ages the age distribution is complex and likely non Gaussian. From the empirical sequences, we find that Gl 410 is slightly younger than Praesepe that is somewhat unlikely to be as young as the Pleiades (120 Myr, although the probability is not 0%), and very unlikely to be as young as the Tucana-Horologium association (45 Myr). Based on these considerations, we propose an age of 480150 Myr for Gl 410. Further dedicated work it is necessary to determine the age of Gl 410 with a higher precision, but this study is beyond the scope of this paper. A possible way to narrow-down the uncertainty, would be to map the distribution of reference age populations at this spectral type, reverse the distribution using Bayes theorem and Gl 410’s rotational period and then obtain a non-gaussian age probability distribution for Gl 410.
A young age for Gl 410 is of interest because it suggest that Gl 410 might be related to the Ursa Major (UMA) cluster which is approximately 400 Myr old. Gl 410 indeed has been previously suggested to be member of UMA by Ammler-von Eiff & Guenther (2009), member of a diffuse region associated with UMA by King et al. (2003) and member of the ”UMA moving group” by Montes et al. (2001). The Galactic velocities of Gl 410 appear consistent with the proposed Ursa Major corona (Gagné et al. 2020).
Appendix B Radial velocity measurements
| DATE | RJD | RV | RV | ||
| [yyyy-mm-dd] | [days] | [km/s] | [km/s] | [m/s] | [m/s] |
| 2010-03-14 | 55269.43680 | -13.8604 | 0.0017 | -21.9 | 1.7 |
| 2010-03-15 | 55270.60680 | -13.8701 | 0.0026 | -31.6 | 2.6 |
| 2010-03-16 | 55271.58010 | -13.8473 | 0.0027 | -8.8 | 2.7 |
| 2010-03-17 | 55272.48330 | -13.8439 | 0.0016 | -5.4 | 1.6 |
| 2010-03-18 | 55273.57290 | -13.8385 | 0.0027 | 0.0 | 2.7 |
| 2010-03-25 | 55280.61610 | -13.8384 | 0.0032 | 0.1 | 3.2 |
| 2010-03-27 | 55282.44940 | -13.8448 | 0.0017 | -6.3 | 1.7 |
| 2010-03-28 | 55283.46120 | -13.8541 | 0.0017 | -15.6 | 1.7 |
| 2010-03-29 | 55284.38640 | -13.8149 | 0.0021 | 23.6 | 2.1 |
| 2010-03-30 | 55285.52760 | -13.8407 | 0.0035 | -2.2 | 3.5 |
| 2010-03-31 | 55286.43530 | -13.8225 | 0.0020 | 16.0 | 2.0 |
| 2010-04-01 | 55287.34720 | -13.8244 | 0.0020 | 14.1 | 2.0 |
| 2010-04-02 | 55288.45810 | -13.8388 | 0.0022 | -0.3 | 2.2 |
| 2010-04-05 | 55291.44930 | -13.8341 | 0.0016 | 4.4 | 1.6 |
| 2010-04-06 | 55292.42410 | -13.8455 | 0.0016 | -7.0 | 1.6 |
| 2010-04-09 | 55295.44570 | -13.8486 | 0.0019 | -10.1 | 1.9 |
| 2010-04-10 | 55296.45340 | -13.8696 | 0.0028 | -31.1 | 2.8 |
| 2010-04-11 | 55297.49250 | -13.8793 | 0.0025 | -40.8 | 2.5 |
| 2010-04-15 | 55301.39210 | -13.8129 | 0.0016 | 25.6 | 1.6 |
| 2010-04-17 | 55303.35330 | -13.8627 | 0.0025 | -24.2 | 2.5 |
| 2010-04-18 | 55304.38490 | -13.8715 | 0.0017 | -33.0 | 1.7 |
| 2011-11-10 | 55875.68740 | -13.8371 | 0.0024 | 1.4 | 2.4 |
| 2012-04-30 | 56047.43600 | -13.8359 | 0.0036 | 2.6 | 3.6 |
| 2012-05-01 | 56048.48910 | -13.8259 | 0.0030 | 12.6 | 3.0 |
| 2012-05-02 | 56049.40220 | -13.8208 | 0.0027 | 17.7 | 2.7 |
| 2012-05-03 | 56050.43920 | -13.8458 | 0.0022 | -7.3 | 2.2 |
| 2012-05-04 | 56051.38670 | -13.8524 | 0.0023 | -13.9 | 2.3 |
| 2012-05-06 | 56053.37910 | -13.8506 | 0.0028 | -12.1 | 2.8 |
| 2012-05-08 | 56055.36830 | -13.8353 | 0.0026 | 3.2 | 2.6 |
| 2012-05-10 | 56057.37370 | -13.8354 | 0.0021 | 3.1 | 2.1 |
| 2012-05-11 | 56058.37440 | -13.8399 | 0.0034 | -1.4 | 3.4 |
| 2012-05-12 | 56059.35330 | -13.8385 | 0.0021 | 0.0 | 2.1 |
| 2012-05-14 | 56061.35010 | -13.8216 | 0.0034 | 16.9 | 3.4 |
| 2012-05-15 | 56062.36800 | -13.8354 | 0.0041 | 3.1 | 4.1 |
| 2012-05-16 | 56063.34620 | -13.8368 | 0.0026 | 1.7 | 2.6 |
| 2012-05-17 | 56064.37340 | -13.8438 | 0.0033 | -5.3 | 3.3 |
| 2012-05-19 | 56066.36940 | -13.8435 | 0.0028 | -5.0 | 2.8 |
| 2012-05-23 | 56070.35420 | -13.8433 | 0.0025 | -4.8 | 2.5 |
| 2012-05-29 | 56076.36430 | -13.8322 | 0.0024 | 6.3 | 2.4 |
| 2012-06-03 | 56081.35990 | -13.8404 | 0.0020 | -1.9 | 2.0 |
| 2012-06-05 | 56083.36530 | -13.8443 | 0.0029 | -5.8 | 2.9 |
| 2012-06-06 | 56084.36400 | -13.8286 | 0.0019 | 9.9 | 1.9 |
| 2012-06-09 | 56087.38290 | -13.8641 | 0.0044 | -25.6 | 4.4 |
| 2012-06-13 | 56091.37860 | -13.8379 | 0.0033 | 0.6 | 3.3 |
| 2012-06-14 | 56092.36600 | -13.8330 | 0.0020 | 5.5 | 2.0 |
| 2021-01-26 | 59240.47760 | -13.8204 | 0.0032 | 18.1 | 3.2 |
| 2021-02-02 | 59247.67160 | -13.8297 | 0.0023 | 8.8 | 2.3 |
| 2021-02-19 | 59264.58910 | -13.8444 | 0.0050 | -5.9 | 5.0 |
| 2021-02-20 | 59265.61360 | -13.8178 | 0.0023 | 20.7 | 2.3 |
| 2021-02-21 | 59266.45130 | -13.8086 | 0.0021 | 29.9 | 2.1 |
| 2021-02-25 | 59270.47650 | -13.8437 | 0.0020 | -5.2 | 2.0 |
| 2021-02-26 | 59271.53270 | -13.8391 | 0.0017 | -0.6 | 1.7 |
| 2021-02-28 | 59273.61860 | -13.8127 | 0.0030 | 25.8 | 3.0 |
| 2021-03-04 | 59277.62460 | -13.8424 | 0.0043 | -3.9 | 4.3 |
| 2021-03-25 | 59298.45260 | -13.8338 | 0.0020 | 4.7 | 2.0 |
| 2021-03-29 | 59302.49460 | -13.8338 | 0.0021 | 4.7 | 2.1 |
| 2021-03-30 | 59303.39620 | -13.8390 | 0.0017 | -0.5 | 1.7 |
| 2021-04-01 | 59305.49890 | -13.8423 | 0.0022 | -3.8 | 2.2 |
| 2021-05-09 | 59343.40690 | -13.8374 | 0.0023 | 1.1 | 2.3 |
| 2021-05-19 | 59353.44370 | -13.8506 | 0.0026 | -12.1 | 2.6 |
| 2021-05-23 | 59357.40020 | -13.8269 | 0.0022 | 11.6 | 2.2 |
| 2021-05-28 | 59362.37750 | -13.8239 | 0.0023 | 14.6 | 2.3 |
| 2021-06-13 | 59378.36710 | -13.8283 | 0.0021 | 10.2 | 2.1 |
| 2021-06-16 | 59381.39670 | -13.8432 | 0.0029 | -4.7 | 2.9 |
| DATE | RJD | RV | RV | ||
| [yyyy-mm-dd] | [days] | [km/s] | [km/s] | [m/s] | [m/s] |
| 2021-06-17 | 59382.36870 | -13.8342 | 0.0033 | 4.3 | 3.3 |
| 2021-12-12 | 59560.68390 | -13.8444 | 0.0021 | -5.9 | 2.1 |
| 2021-12-16 | 59564.60700 | -13.8241 | 0.0018 | 14.4 | 1.8 |
| 2021-12-18 | 59566.66560 | -13.8539 | 0.0018 | -15.4 | 1.8 |
| 2021-12-20 | 59568.62540 | -13.8302 | 0.0018 | 8.3 | 1.8 |
| 2022-01-14 | 59593.58150 | -13.8458 | 0.0023 | -7.3 | 2.3 |
| 2022-01-16 | 59595.61930 | -13.8516 | 0.0019 | -13.1 | 1.9 |
| 2022-01-18 | 59597.62530 | -13.8274 | 0.0020 | 11.1 | 2.0 |
| 2022-01-24 | 59603.66480 | -13.8323 | 0.0024 | 6.2 | 2.4 |
| 2022-01-27 | 59606.50480 | -13.8318 | 0.0024 | 6.7 | 2.4 |
| 2022-01-31 | 59610.58680 | -13.8390 | 0.0028 | -0.5 | 2.8 |
| 2022-02-11 | 59621.62300 | -13.8383 | 0.0032 | 0.2 | 3.2 |
| 2022-02-15 | 59625.66100 | -13.8438 | 0.0026 | -5.3 | 2.6 |
| 2022-02-18 | 59628.49180 | -13.8260 | 0.0033 | 12.5 | 3.3 |
| 2021-03-04 | 59277.62460 | -13.8424 | 0.0043 | -3.9 | 4.3 |
| 2021-03-25 | 59298.45260 | -13.8338 | 0.0020 | 4.7 | 2.0 |
| 2021-03-29 | 59302.49460 | -13.8338 | 0.0021 | 4.7 | 2.1 |
| 2021-03-30 | 59303.39620 | -13.8390 | 0.0017 | -0.5 | 1.7 |
| 2021-04-01 | 59305.49890 | -13.8423 | 0.0022 | -3.8 | 2.2 |
| 2021-05-09 | 59343.40690 | -13.8374 | 0.0023 | 1.1 | 2.3 |
| 2021-05-19 | 59353.44370 | -13.8506 | 0.0026 | -12.1 | 2.6 |
| 2021-05-23 | 59357.40020 | -13.8269 | 0.0022 | 11.6 | 2.2 |
| 2021-05-28 | 59362.37750 | -13.8239 | 0.0023 | 14.6 | 2.3 |
| 2021-06-13 | 59378.36710 | -13.8283 | 0.0021 | 10.2 | 2.1 |
| 2021-06-16 | 59381.39670 | -13.8432 | 0.0029 | -4.7 | 2.9 |
| 2021-06-17 | 59382.36870 | -13.8342 | 0.0033 | 4.3 | 3.3 |
| 2021-12-12 | 59560.68390 | -13.8444 | 0.0021 | -5.9 | 2.1 |
| 2021-12-18 | 59566.66560 | -13.8539 | 0.0018 | -15.4 | 1.8 |
| 2021-12-20 | 59568.62540 | -13.8302 | 0.0018 | 8.3 | 1.8 |
| 2022-01-14 | 59593.58150 | -13.8458 | 0.0023 | -7.3 | 2.3 |
| 2022-01-16 | 59595.61930 | -13.8516 | 0.0019 | -13.1 | 1.9 |
| 2022-01-18 | 59597.62530 | -13.8274 | 0.0020 | 11.1 | 2.0 |
| 2022-01-24 | 59603.66480 | -13.8323 | 0.0024 | 6.2 | 2.4 |
| 2022-01-27 | 59606.50480 | -13.8318 | 0.0024 | 6.7 | 2.4 |
| 2022-01-31 | 59610.58680 | -13.8390 | 0.0028 | -0.5 | 2.8 |
| 2022-02-11 | 59621.62300 | -13.8383 | 0.0032 | 0.2 | 3.2 |
| 2022-02-15 | 59625.66100 | -13.8438 | 0.0026 | -5.3 | 2.6 |
| 2022-02-18 | 59628.49180 | -13.8260 | 0.0033 | 12.5 | 3.3 |
| 2022-02-23 | 59633.45240 | -13.8138 | 0.0028 | 24.7 | 2.8 |
| 2022-02-27 | 59637.61630 | -13.8606 | 0.0026 | -22.1 | 2.6 |
| 2022-03-05 | 59643.48510 | -13.8510 | 0.0028 | -12.5 | 2.8 |
| 2022-03-23 | 59661.47130 | -13.8335 | 0.0022 | 5.0 | 2.2 |
| 2022-04-11 | 59680.43830 | -13.8567 | 0.0028 | -18.2 | 2.8 |
| 2022-04-17 | 59686.51640 | -13.8512 | 0.0023 | -12.7 | 2.3 |
| 2022-05-02 | 59701.42880 | -13.8246 | 0.0027 | 13.9 | 2.7 |
| 2022-05-07 | 59706.47180 | -13.8437 | 0.0037 | -5.2 | 3.7 |
| 2022-05-14 | 59713.38900 | -13.8398 | 0.0025 | -1.3 | 2.5 |
| 2022-05-18 | 59717.37400 | -13.8277 | 0.0027 | 10.8 | 2.7 |
| 2022-05-22 | 59721.42130 | -13.8475 | 0.0031 | -9.0 | 3.1 |
| 2022-05-29 | 59728.40620 | -13.8409 | 0.0034 | -2.4 | 3.4 |
| 2022-06-06 | 59736.40040 | -13.8239 | 0.0026 | 14.6 | 2.6 |
| 2022-06-11 | 59741.40070 | -13.8353 | 0.0029 | 3.2 | 2.9 |
| 2022-06-14 | 59744.37810 | -13.8167 | 0.0032 | 21.8 | 3.2 |
| 2022-12-07 | 59920.71050 | -13.8534 | 0.0026 | -14.9 | 2.6 |
| 2022-12-08 | 59921.61120 | -13.8226 | 0.0022 | 15.9 | 2.2 |
| 2023-01-11 | 59955.57580 | -13.8338 | 0.0026 | 4.7 | 2.6 |
| 2023-01-29 | 59973.55860 | -13.8611 | 0.0024 | -22.6 | 2.4 |
| 2023-02-01 | 59976.55640 | -13.8468 | 0.0027 | -8.3 | 2.7 |
| 2023-02-02 | 59977.61200 | -13.8297 | 0.0039 | 8.8 | 3.9 |
| 2023-02-08 | 59983.51960 | -13.8146 | 0.0024 | 23.9 | 2.4 |
| 2023-02-13 | 59988.50910 | -13.8575 | 0.0019 | -19.0 | 1.9 |
| 2023-02-15 | 59990.59050 | -13.8330 | 0.0019 | 5.5 | 1.9 |
| 2023-03-04 | 60007.46200 | -13.8173 | 0.0021 | 21.2 | 2.1 |
| LBL | Wapiti | |||||||||
| DATE | RJD | RV | RV | RV | RV | |||||
| [yyyy-mm-dd] | [days] | [km/s] | [km/s] | [km/s] | [m/s] | [m/s] | [km/s] | [km/s] | [m/s] | [m/s] |
| 2020-11-04 | 59158.11239 | -40.58 | -13.8818 | 0.0023 | 15.88 | 2.3 | -13.9924 | 0.0030 | 11.00 | 3.0 |
| 2020-11-05 | 59159.10674 | -40.80 | -13.8875 | 0.0023 | 10.22 | 2.3 | -14.0006 | 0.0030 | 2.83 | 3.0 |
| 2020-12-24 | 59208.05450 | -40.68 | -13.8916 | 0.0028 | 6.16 | 2.8 | -13.9993 | 0.0036 | 4.10 | 3.6 |
| 2020-12-25 | 59209.10708 | -40.32 | -13.8985 | 0.0023 | -0.80 | 2.3 | -14.0033 | 0.0030 | 0.12 | 3.0 |
| 2020-12-26 | 59210.04378 | -40.25 | -13.9089 | 0.0023 | -11.19 | 2.3 | -14.0078 | 0.0030 | -4.40 | 3.0 |
| 2020-12-29 | 59213.12163 | -39.29 | -13.8912 | 0.0027 | 6.50 | 2.7 | -14.0024 | 0.0034 | 0.99 | 3.4 |
| 2020-12-30 | 59214.12312 | -39.03 | -13.8926 | 0.0023 | 5.12 | 2.3 | -14.0031 | 0.0029 | 0.28 | 2.9 |
| 2020-12-31 | 59215.09706 | -38.83 | -13.8988 | 0.0022 | -1.08 | 2.2 | -14.0015 | 0.0029 | 1.94 | 2.9 |
| 2021-01-02 | 59217.14901 | -38.11 | -13.8899 | 0.0021 | 7.87 | 2.1 | -14.0018 | 0.0027 | 1.62 | 2.7 |
| 2021-01-03 | 59218.07681 | -38.02 | -13.8854 | 0.0022 | 12.29 | 2.2 | -14.0033 | 0.0030 | 0.15 | 3.0 |
| 2021-01-04 | 59219.05500 | -37.78 | -13.8915 | 0.0024 | 6.25 | 2.4 | -14.0084 | 0.0031 | -4.96 | 3.1 |
| 2021-01-05 | 59220.04821 | -37.49 | -13.8914 | 0.0022 | 6.35 | 2.2 | -13.9996 | 0.0029 | 3.80 | 2.9 |
| 2021-01-06 | 59221.08802 | -37.07 | -13.8944 | 0.0023 | 3.34 | 2.3 | -14.0060 | 0.0030 | -2.54 | 3.0 |
| 2021-01-07 | 59222.09372 | -36.73 | -13.8970 | 0.0021 | 0.72 | 2.1 | -14.0081 | 0.0028 | -4.66 | 2.8 |
| 2021-01-08 | 59223.10069 | -36.37 | -13.9014 | 0.0021 | -3.65 | 2.1 | -14.0060 | 0.0028 | -2.56 | 2.8 |
| 2021-02-21 | 59266.98306 | -17.01 | -13.9008 | 0.0025 | -3.09 | 2.5 | -13.9984 | 0.0031 | 5.05 | 3.1 |
| 2021-02-22 | 59268.03968 | -16.33 | -13.8913 | 0.0022 | 6.45 | 2.2 | -14.0012 | 0.0029 | 2.18 | 2.9 |
| 2021-02-23 | 59269.12577 | -15.61 | -13.8943 | 0.0025 | 3.47 | 2.5 | - | - | - | - |
| 2021-02-26 | 59271.93626 | -14.60 | -13.8951 | 0.0021 | 2.60 | 2.1 | -13.9993 | 0.0027 | 4.17 | 2.7 |
| 2021-02-28 | 59274.11250 | -13.10 | -13.8880 | 0.0021 | 9.68 | 2.1 | -14.0053 | 0.0028 | -1.88 | 2.8 |
| 2021-03-02 | 59276.09380 | -12.12 | -13.8996 | 0.0023 | -1.89 | 2.3 | -14.0085 | 0.0030 | -5.08 | 3.0 |
| 2021-03-03 | 59277.01475 | -11.82 | -13.8925 | 0.0022 | 5.22 | 2.2 | -14.0062 | 0.0028 | -2.76 | 2.8 |
| 2021-03-04 | 59278.07948 | -11.14 | -13.8974 | 0.0022 | 0.32 | 2.2 | -14.0082 | 0.0029 | -4.78 | 2.9 |
| 2021-03-20 | 59294.04638 | -3.34 | -13.9233 | 0.0021 | -25.55 | 2.1 | -14.0066 | 0.0028 | -3.17 | 2.8 |
| 2021-03-21 | 59295.00581 | -2.96 | -13.9073 | 0.0022 | -9.61 | 2.2 | -13.9990 | 0.0030 | 4.38 | 3.0 |
| 2021-03-22 | 59296.06857 | -2.37 | -13.9263 | 0.0024 | -28.61 | 2.4 | -14.0116 | 0.0030 | -8.16 | 3.0 |
| 2021-03-23 | 59297.00310 | -2.04 | -13.9027 | 0.0022 | -4.95 | 2.2 | -13.9927 | 0.0029 | 10.76 | 2.9 |
| 2021-03-24 | 59298.05160 | -1.49 | -13.9072 | 0.0022 | -9.43 | 2.2 | -14.0061 | 0.0029 | -2.67 | 2.9 |
| 2021-03-26 | 59300.02528 | -0.64 | -13.9151 | 0.0029 | -17.35 | 2.9 | -14.0013 | 0.0036 | 2.11 | 3.6 |
| 2021-03-27 | 59300.80238 | -0.81 | -13.9069 | 0.0022 | -9.17 | 2.2 | -14.0021 | 0.0029 | 1.28 | 2.9 |
| 2021-03-31 | 59304.97415 | 1.41 | -13.8805 | 0.0023 | 17.23 | 2.3 | -13.9918 | 0.0030 | 11.62 | 3.0 |
| 2021-04-01 | 59306.02154 | 1.93 | -13.8782 | 0.0022 | 19.57 | 2.2 | -13.9856 | 0.0029 | 17.84 | 2.9 |
| 2021-04-22 | 59326.95738 | 9.36 | -13.9010 | 0.0023 | -3.32 | 2.3 | -14.0053 | 0.0030 | -1.93 | 3.0 |
| 2021-04-23 | 59327.87646 | 9.46 | -13.8901 | 0.0022 | 7.62 | 2.2 | -13.9992 | 0.0028 | 4.18 | 2.8 |
| 2021-04-24 | 59328.96195 | 9.93 | -13.8936 | 0.0022 | 4.16 | 2.2 | -14.0015 | 0.0029 | 1.88 | 2.9 |
| 2021-04-25 | 59329.91740 | 10.12 | -13.9063 | 0.0027 | -8.52 | 2.7 | -14.0040 | 0.0034 | -0.56 | 3.4 |
| 2021-04-26 | 59330.94314 | 10.43 | -13.8948 | 0.0022 | 2.92 | 2.2 | - | - | - | - |
| 2021-04-27 | 59331.96518 | 10.73 | -13.8938 | 0.0023 | 3.96 | 2.3 | -14.0026 | 0.0031 | 0.82 | 3.1 |
| 2021-04-28 | 59332.96320 | 10.98 | -13.8992 | 0.0022 | -1.42 | 2.2 | -14.0065 | 0.0028 | -3.13 | 2.8 |
| 2021-04-30 | 59334.93467 | 11.42 | -13.9009 | 0.0022 | -3.19 | 2.2 | -14.0056 | 0.0029 | -2.17 | 2.9 |
| 2021-05-01 | 59335.94049 | 11.66 | -13.8995 | 0.0021 | -1.73 | 2.1 | -14.0067 | 0.0027 | -3.26 | 2.7 |
| 2021-05-02 | 59336.93193 | 11.87 | -13.9005 | 0.0023 | -2.78 | 2.3 | -14.0072 | 0.0029 | -3.78 | 2.9 |
| 2021-05-03 | 59337.91950 | 12.07 | -13.9042 | 0.0024 | -6.42 | 2.4 | -14.0097 | 0.0031 | -6.25 | 3.1 |
| 2021-06-19 | 59384.77577 | 13.12 | -13.8991 | 0.0021 | -1.42 | 2.1 | -14.0077 | 0.0026 | -4.31 | 2.6 |
| 2021-06-20 | 59385.77368 | 12.95 | -13.8935 | 0.0023 | 4.22 | 2.3 | -14.0001 | 0.0029 | 3.29 | 2.9 |
| 2021-06-21 | 59386.76976 | 12.78 | -13.8942 | 0.0022 | 3.58 | 2.2 | -14.0030 | 0.0029 | 0.43 | 2.9 |
| 2021-06-23 | 59388.75094 | 12.38 | -13.8874 | 0.0024 | 10.35 | 2.4 | -14.0001 | 0.0031 | 3.32 | 3.1 |
| 2021-06-24 | 59389.74885 | 12.19 | -13.9032 | 0.0023 | -5.52 | 2.3 | -14.0065 | 0.0030 | -3.03 | 3.0 |
| 2021-06-25 | 59390.75487 | 12.01 | -13.8926 | 0.0022 | 5.14 | 2.2 | -13.9966 | 0.0028 | 6.79 | 2.8 |
| 2021-06-26 | 59391.74876 | 11.79 | -13.8989 | 0.0022 | -1.15 | 2.2 | -13.9998 | 0.0028 | 3.60 | 2.8 |
| 2021-06-27 | 59392.75010 | 11.59 | -13.8984 | 0.0022 | -0.63 | 2.2 | -14.0063 | 0.0030 | -2.86 | 3.0 |
| 2021-06-28 | 59393.75228 | 11.37 | -13.8907 | 0.0024 | 6.99 | 2.4 | -13.9965 | 0.0032 | 6.89 | 3.2 |
| 2021-06-29 | 59394.74804 | 11.14 | -13.8978 | 0.0024 | -0.02 | 2.4 | -14.0013 | 0.0032 | 2.12 | 3.2 |
| 2021-06-30 | 59395.74896 | 10.92 | -13.9049 | 0.0026 | -7.21 | 2.6 | -14.0069 | 0.0033 | -3.47 | 3.3 |
| 2021-07-01 | 59396.74897 | 10.68 | -13.9084 | 0.0023 | -10.71 | 2.3 | -14.0099 | 0.0030 | -6.49 | 3.0 |
| 2021-07-02 | 59397.74861 | 10.43 | -13.9082 | 0.0029 | -10.51 | 2.9 | -14.0091 | 0.0036 | -5.68 | 3.6 |
| 2021-07-17 | 59412.74636 | 5.93 | -13.9049 | 0.0022 | -7.18 | 2.2 | -14.0080 | 0.0028 | -4.56 | 2.8 |
| 2021-07-19 | 59414.73953 | 5.22 | -13.9142 | 0.0024 | -16.46 | 2.4 | -14.0114 | 0.0031 | -7.98 | 3.1 |
| 2021-11-16 | 59535.13288 | -42.51 | -13.8979 | 0.0021 | -0.19 | 2.1 | -14.0048 | 0.0028 | -1.35 | 2.8 |
| 2021-11-17 | 59536.15385 | -42.58 | -13.9002 | 0.0022 | -2.50 | 2.2 | -14.0066 | 0.0029 | -3.16 | 2.9 |
| 2021-11-18 | 59537.12805 | -42.73 | -13.8940 | 0.0021 | 3.74 | 2.1 | -14.0013 | 0.0029 | 2.15 | 2.9 |
| 2021-11-19 | 59538.15074 | -42.78 | -13.8831 | 0.0023 | 14.63 | 2.3 | -13.9934 | 0.0030 | 10.07 | 3.0 |
| 2021-11-20 | 59539.15254 | -42.86 | -13.8931 | 0.0015 | 4.63 | 1.5 | -14.0046 | 0.0019 | -1.15 | 1.9 |
| 2021-11-21 | 59540.13474 | -42.98 | -13.8994 | 0.0020 | -1.69 | 2.0 | -14.0084 | 0.0027 | -5.00 | 2.7 |
| 2021-11-22 | 59541.12292 | -43.07 | -13.8914 | 0.0023 | 6.30 | 2.3 | -14.0006 | 0.0029 | 2.85 | 2.9 |
| 2021-12-09 | 59558.14083 | -42.84 | -13.8931 | 0.0021 | 4.61 | 2.1 | -14.0016 | 0.0028 | 1.81 | 2.8 |
| 2021-12-10 | 59559.12174 | -42.79 | -13.8975 | 0.0024 | 0.25 | 2.4 | -14.0067 | 0.0031 | -3.26 | 3.1 |
| 2021-12-11 | 59560.16381 | -42.58 | -13.8946 | 0.0023 | 3.12 | 2.3 | -14.0068 | 0.0030 | -3.42 | 3.0 |
| 2021-12-13 | 59562.12940 | -42.43 | -13.8839 | 0.0021 | 13.81 | 2.1 | -13.9955 | 0.0028 | 7.95 | 2.8 |
| 2021-12-15 | 59564.11415 | -42.20 | -13.8964 | 0.0021 | 1.32 | 2.1 | -14.0015 | 0.0028 | 1.94 | 2.8 |
| 2022-03-12 | 59650.96249 | -7.54 | -13.8859 | 0.0022 | 11.87 | 2.2 | -14.0022 | 0.0028 | 1.19 | 2.8 |
| 2022-03-13 | 59651.98257 | -6.99 | -13.8748 | 0.0022 | 22.93 | 2.2 | -13.9936 | 0.0029 | 9.86 | 2.9 |
| 2022-03-14 | 59652.86144 | -6.86 | -13.8788 | 0.0022 | 18.94 | 2.2 | -14.0000 | 0.0029 | 3.46 | 2.9 |
| 2022-03-15 | 59654.02433 | -5.90 | -13.8797 | 0.0022 | 18.06 | 2.2 | -13.9950 | 0.0028 | 8.38 | 2.8 |
| 2022-03-16 | 59655.06137 | -5.33 | -13.8836 | 0.0022 | 14.14 | 2.2 | -13.9976 | 0.0027 | 5.80 | 2.7 |
| 2022-03-17 | 59655.97333 | -5.07 | -13.8894 | 0.0022 | 8.31 | 2.2 | -14.0060 | 0.0029 | -2.54 | 2.9 |
| 2022-03-18 | 59657.06403 | -4.38 | -13.8991 | 0.0025 | -1.41 | 2.5 | -14.0006 | 0.0032 | 2.79 | 3.2 |
| 2022-03-19 | 59658.05424 | -3.92 | -13.9047 | 0.0021 | -7.00 | 2.1 | -14.0051 | 0.0028 | -1.66 | 2.8 |
| 2022-03-20 | 59658.99954 | -3.58 | -13.9136 | 0.0021 | -15.83 | 2.1 | -14.0093 | 0.0028 | -5.86 | 2.8 |
| 2022-03-22 | 59660.99607 | -2.65 | -13.9202 | 0.0022 | -22.44 | 2.2 | -14.0069 | 0.0028 | -3.46 | 2.8 |
| 2022-03-23 | 59661.99049 | -2.21 | -13.9218 | 0.0018 | -24.02 | 1.8 | - | - | - | - |
| 2022-04-09 | 59678.99154 | 4.99 | -13.9011 | 0.0024 | -3.36 | 2.4 | -14.0100 | 0.0028 | -6.63 | 2.8 |
| 2022-04-12 | 59681.97934 | 6.06 | -13.8998 | 0.0023 | -2.07 | 2.3 | -14.0067 | 0.0030 | -3.29 | 3.0 |
| 2022-04-13 | 59682.99757 | 6.45 | -13.8968 | 0.0022 | 0.95 | 2.2 | -14.0048 | 0.0028 | -1.42 | 2.8 |
| 2022-04-14 | 59684.01154 | 6.81 | -13.8981 | 0.0031 | -0.38 | 3.1 | -14.0043 | 0.0037 | -0.93 | 3.7 |
| 2022-04-15 | 59684.98764 | 7.12 | -13.8997 | 0.0023 | -1.97 | 2.3 | -14.0130 | 0.0030 | -9.58 | 3.0 |
| 2022-04-18 | 59687.98838 | 8.10 | -13.8923 | 0.0023 | 5.47 | 2.3 | -14.0015 | 0.0031 | 1.93 | 3.1 |
| 2022-04-19 | 59688.96958 | 8.39 | -13.8895 | 0.0023 | 8.26 | 2.3 | -14.0002 | 0.0031 | 3.25 | 3.1 |
| 2022-04-20 | 59689.95600 | 8.67 | -13.8948 | 0.0026 | 2.90 | 2.6 | -13.9974 | 0.0034 | 6.06 | 3.4 |
| 2022-04-21 | 59690.95481 | 8.98 | -13.9014 | 0.0030 | -3.69 | 3.0 | - | - | - | - |
| 2022-05-11 | 59710.90469 | 13.47 | -13.8952 | 0.0023 | 2.51 | 2.3 | -14.0079 | 0.0030 | -4.47 | 3.0 |
| 2022-05-12 | 59711.91277 | 13.63 | -13.8911 | 0.0022 | 6.63 | 2.2 | -14.0012 | 0.0029 | 2.27 | 2.9 |
| 2022-05-13 | 59712.88788 | 13.73 | -13.8934 | 0.0022 | 4.33 | 2.2 | -14.0003 | 0.0029 | 3.14 | 2.9 |
| 2022-05-16 | 59715.85928 | 14.05 | -13.8964 | 0.0022 | 1.34 | 2.2 | -14.0063 | 0.0029 | -2.87 | 2.9 |
| 2022-05-17 | 59716.86406 | 14.17 | -13.9011 | 0.0024 | -3.40 | 2.4 | -14.0059 | 0.0031 | -2.47 | 3.1 |
| 2022-06-02 | 59732.77623 | 14.69 | -13.9027 | 0.0022 | -4.95 | 2.2 | -14.0147 | 0.0028 | -11.31 | 2.8 |
| 2022-06-03 | 59733.77334 | 14.66 | -13.9098 | 0.0023 | -12.06 | 2.3 | -14.0164 | 0.0030 | -13.02 | 3.0 |
| 2022-06-05 | 59735.77903 | 14.59 | -13.8974 | 0.0022 | 0.35 | 2.2 | -14.0069 | 0.0028 | -3.52 | 2.8 |
| 2022-06-06 | 59736.84702 | 14.66 | -13.8861 | 0.0022 | 11.62 | 2.2 | -13.9974 | 0.0028 | 6.04 | 2.8 |
| 2022-06-07 | 59737.85329 | 14.60 | -13.8896 | 0.0022 | 8.10 | 2.2 | - | - | - | - |
| 2022-06-08 | 59738.75236 | 14.36 | -13.8964 | 0.0021 | 1.37 | 2.1 | -14.0034 | 0.0027 | 0.00 | 2.7 |
| 2022-06-09 | 59739.75308 | 14.28 | -13.8951 | 0.0022 | 2.59 | 2.2 | -14.0020 | 0.0029 | 1.42 | 2.9 |
| 2022-06-10 | 59740.79830 | 14.29 | -13.9005 | 0.0022 | -2.81 | 2.2 | -14.0087 | 0.0029 | -5.30 | 2.9 |
| 2022-06-11 | 59741.78115 | 14.17 | -13.8851 | 0.0023 | 12.59 | 2.3 | -13.9990 | 0.0030 | 4.43 | 3.0 |
| 2022-06-12 | 59742.79232 | 14.09 | -13.8813 | 0.0028 | 16.46 | 2.8 | -13.9974 | 0.0034 | 6.05 | 3.4 |
| 2022-06-13 | 59743.75821 | 13.92 | -13.8899 | 0.0032 | 7.86 | 3.2 | -14.0024 | 0.0039 | 1.07 | 3.9 |
| 2022-11-06 | 59890.13221 | -40.85 | -13.8987 | 0.0022 | -1.00 | 2.2 | -13.9998 | 0.0029 | 3.66 | 2.9 |
| 2022-11-09 | 59893.12159 | -41.44 | -13.8974 | 0.0020 | 0.28 | 2.0 | -14.0021 | 0.0026 | 1.31 | 2.6 |
| 2022-11-12 | 59896.11941 | -41.94 | -13.8999 | 0.0022 | -2.13 | 2.2 | -14.0020 | 0.0030 | 1.46 | 3.0 |
| 2022-11-15 | 59899.15299 | -42.31 | -13.8978 | 0.0021 | -0.08 | 2.1 | -13.9979 | 0.0027 | 5.48 | 2.7 |
| 2022-11-18 | 59902.15197 | -42.66 | -13.9045 | 0.0022 | -6.81 | 2.2 | -14.0025 | 0.0029 | 0.93 | 2.9 |
| 2022-11-19 | 59903.15749 | -42.75 | -13.8981 | 0.0023 | -0.35 | 2.3 | -14.0010 | 0.0028 | 2.42 | 2.8 |
| 2022-12-30 | 59944.16181 | -39.04 | -13.9027 | 0.0023 | -5.00 | 2.3 | -13.9983 | 0.0030 | 5.13 | 3.0 |
| 2022-12-31 | 59945.15901 | -38.78 | -13.9060 | 0.0023 | -8.29 | 2.3 | -14.0067 | 0.0029 | -3.31 | 2.9 |
| 2023-01-01 | 59946.15454 | -38.51 | -13.8977 | 0.0022 | 0.02 | 2.2 | -14.0036 | 0.0028 | -0.13 | 2.8 |
| 2023-01-04 | 59949.14985 | -37.64 | -13.8898 | 0.0023 | 7.95 | 2.3 | -13.9953 | 0.0030 | 8.17 | 3.0 |
| 2023-01-05 | 59950.17815 | -37.26 | -13.9077 | 0.0022 | -10.00 | 2.2 | -13.9998 | 0.0029 | 3.58 | 2.9 |
| 2023-01-06 | 59951.14762 | -37.03 | -13.9084 | 0.0020 | -10.63 | 2.0 | -14.0044 | 0.0027 | -1.01 | 2.7 |
| 2023-01-08 | 59953.18602 | -36.28 | -13.9015 | 0.0020 | -3.81 | 2.0 | -14.0086 | 0.0026 | -5.21 | 2.6 |
| 2023-01-09 | 59954.15923 | -36.01 | -13.8914 | 0.0020 | 6.38 | 2.0 | - | - | - | - |
| 2023-01-11 | 59956.18776 | -35.25 | -13.9068 | 0.0021 | -9.04 | 2.1 | - | - | - | - |
| 2023-01-13 | 59958.12996 | -34.69 | -13.8939 | 0.0022 | 3.84 | 2.2 | - | - | - | - |
| 2023-02-02 | 59978.09021 | -26.34 | -13.8842 | 0.0019 | 13.57 | 1.9 | - | - | - | - |
| 2023-02-08 | 59984.11968 | -23.40 | -13.8840 | 0.0023 | 13.69 | 2.3 | - | - | - | - |
| 2023-03-31 | 60035.00130 | 1.28 | -13.8792 | 0.0022 | 18.56 | 2.2 | - | - | - | - |
| 2023-04-07 | 60041.95876 | 4.03 | -13.8727 | 0.0022 | 25.00 | 2.2 | - | - | - | - |
| 2023-04-11 | 60045.99173 | 5.62 | -13.8981 | 0.0022 | -0.38 | 2.2 | - | - | - | - |
| 2023-04-29 | 60063.90176 | 11.02 | -13.9008 | 0.0021 | -3.03 | 2.1 | - | - | - | - |
| 2023-05-02 | 60066.90070 | 11.72 | -13.9003 | 0.0025 | -2.57 | 2.5 | - | - | - | - |
| 2023-05-05 | 60069.86886 | 12.28 | -13.8935 | 0.0023 | 4.22 | 2.3 | - | - | - | - |
| 2023-05-09 | 60073.87663 | 13.05 | -13.9027 | 0.0023 | -5.02 | 2.3 | - | - | - | - |
| 2023-06-03 | 60098.77913 | 14.67 | -13.8859 | 0.0022 | 11.81 | 2.2 | - | - | - | - |
| 2023-06-07 | 60102.79658 | 14.53 | -13.9005 | 0.0022 | -2.81 | 2.2 | - | - | - | - |
| 2023-06-28 | 60123.74194 | 11.48 | -13.9016 | 0.0024 | -3.83 | 2.4 | - | - | - | - |
| 2023-07-01 | 60126.73580 | 10.78 | -13.9040 | 0.0021 | -6.30 | 2.1 | - | - | - | - |
| 2023-07-04 | 60129.75396 | 10.06 | -13.9014 | 0.0030 | -3.66 | 3.0 | - | - | - | - |
| 2023-07-17 | 60142.73530 | 6.13 | -13.9240 | 0.0025 | -26.30 | 2.5 | - | - | - | - |
| DATE | RJD | RV | RV | d2v | d3v | dLW | |||||||
| [yyyy-mm-dd] | [days] | [km/s] | [km/s] | [m/s] | [m/s] | [m2/s2] | [m2/s2] | [m3/s3] | [m3/s3] | [m/s] | [m/s] | [K] | [K] |
| 2020-11-04 | 59158.11239 | -13.7958 | 0.0023 | 12.10 | 2.3 | -3.05e+04 | 5.87e+03 | 1.39e+07 | 1.17e+07 | -1.69e+05 | 3.25e+04 | -17.4 | 0.5 |
| 2020-11-05 | 59159.10674 | -13.7984 | 0.0022 | 9.50 | 2.2 | -4.46e+04 | 5.85e+03 | 5.51e+06 | 1.16e+07 | -2.47e+05 | 3.24e+04 | -22.9 | 0.5 |
| 2020-12-24 | 59208.05450 | -13.7967 | 0.0027 | 11.12 | 2.7 | 5.45e+04 | 7.18e+03 | -3.94e+06 | 1.43e+07 | 3.02e+05 | 3.98e+04 | 6.1 | 0.6 |
| 2020-12-25 | 59209.10708 | -13.8064 | 0.0022 | 1.47 | 2.2 | 4.42e+04 | 5.82e+03 | -9.32e+06 | 1.16e+07 | 2.45e+05 | 3.23e+04 | 5.5 | 0.5 |
| 2020-12-26 | 59210.04378 | -13.8147 | 0.0022 | -6.82 | 2.2 | 5.90e+04 | 5.69e+03 | -9.01e+06 | 1.13e+07 | 3.27e+05 | 3.16e+04 | 6.1 | 0.4 |
| 2020-12-29 | 59213.12163 | -13.8023 | 0.0026 | 5.54 | 2.6 | 5.13e+03 | 6.83e+03 | 6.90e+06 | 1.36e+07 | 2.84e+04 | 3.79e+04 | -6.0 | 0.5 |
| 2020-12-30 | 59214.12312 | -13.8037 | 0.0023 | 4.16 | 2.3 | 1.31e+03 | 5.84e+03 | -1.00e+07 | 1.16e+07 | 7.27e+03 | 3.24e+04 | -12.2 | 0.5 |
| 2020-12-31 | 59215.09706 | -13.8041 | 0.0022 | 3.77 | 2.2 | -1.42e+04 | 5.65e+03 | 1.17e+07 | 1.12e+07 | -7.86e+04 | 3.13e+04 | -13.9 | 0.4 |
| 2021-01-02 | 59217.14901 | -13.8011 | 0.0020 | 6.78 | 2.0 | -6.76e+04 | 5.28e+03 | -2.20e+06 | 1.04e+07 | -3.75e+05 | 2.93e+04 | -7.5 | 0.4 |
| 2021-01-03 | 59218.07681 | -13.8007 | 0.0022 | 7.11 | 2.2 | 4.67e+03 | 5.74e+03 | 1.30e+07 | 1.13e+07 | 2.59e+04 | 3.18e+04 | -0.0 | 0.4 |
| 2021-01-04 | 59219.05500 | -13.8058 | 0.0023 | 2.10 | 2.3 | 3.76e+04 | 6.07e+03 | 4.16e+06 | 1.20e+07 | 2.09e+05 | 3.36e+04 | 4.6 | 0.5 |
| 2021-01-05 | 59220.04821 | -13.8049 | 0.0021 | 2.94 | 2.1 | 4.13e+04 | 5.54e+03 | 4.57e+06 | 1.09e+07 | 2.29e+05 | 3.07e+04 | 8.6 | 0.4 |
| 2021-01-06 | 59221.08802 | -13.8086 | 0.0023 | -0.79 | 2.3 | 1.69e+04 | 5.90e+03 | -4.98e+06 | 1.16e+07 | 9.35e+04 | 3.27e+04 | 7.0 | 0.4 |
| 2021-01-07 | 59222.09372 | -13.8097 | 0.0021 | -1.84 | 2.1 | 1.98e+04 | 5.38e+03 | 2.16e+06 | 1.06e+07 | 1.10e+05 | 2.98e+04 | 6.9 | 0.4 |
| 2021-01-08 | 59223.10069 | -13.8099 | 0.0021 | -2.01 | 2.1 | 4.88e+04 | 5.41e+03 | 1.14e+06 | 1.07e+07 | 2.70e+05 | 3.00e+04 | 7.5 | 0.4 |
| 2021-02-21 | 59266.98306 | -13.8120 | 0.0024 | -4.16 | 2.4 | -2.38e+03 | 6.26e+03 | 1.26e+07 | 1.25e+07 | -1.32e+04 | 3.47e+04 | -2.2 | 0.5 |
| 2021-02-22 | 59268.03968 | -13.8005 | 0.0021 | 7.33 | 2.1 | -1.96e+04 | 5.55e+03 | -5.66e+06 | 1.10e+07 | -1.09e+05 | 3.08e+04 | -9.8 | 0.4 |
| 2021-02-23 | 59269.12577 | -13.8004 | 0.0024 | 7.49 | 2.4 | -2.12e+04 | 6.30e+03 | 1.22e+07 | 1.25e+07 | -1.18e+05 | 3.49e+04 | -14.3 | 0.5 |
| 2021-02-26 | 59271.93626 | -13.7961 | 0.0020 | 11.79 | 2.0 | -2.49e+04 | 5.24e+03 | -2.10e+06 | 1.03e+07 | -1.38e+05 | 2.90e+04 | -9.3 | 0.4 |
| 2021-02-28 | 59274.11250 | -13.7964 | 0.0020 | 11.47 | 2.0 | 1.30e+04 | 5.20e+03 | -2.40e+07 | 1.03e+07 | 7.22e+04 | 2.88e+04 | 0.9 | 0.4 |
| 2021-03-02 | 59276.09380 | -13.8097 | 0.0023 | -1.81 | 2.3 | 3.06e+04 | 5.82e+03 | -6.76e+06 | 1.15e+07 | 1.70e+05 | 3.23e+04 | 11.3 | 0.5 |
| 2021-03-03 | 59277.01475 | -13.8021 | 0.0021 | 5.79 | 2.1 | 1.93e+04 | 5.52e+03 | -1.75e+07 | 1.08e+07 | 1.07e+05 | 3.06e+04 | 11.4 | 0.4 |
| 2021-03-04 | 59278.07948 | -13.8092 | 0.0021 | -1.30 | 2.1 | -7.46e+03 | 5.53e+03 | 1.81e+07 | 1.08e+07 | -4.13e+04 | 3.07e+04 | 11.1 | 0.4 |
| 2021-03-20 | 59294.04638 | -13.8113 | 0.0020 | -3.40 | 2.0 | 2.68e+04 | 5.28e+03 | 1.53e+07 | 1.05e+07 | 1.49e+05 | 2.93e+04 | 3.3 | 0.4 |
| 2021-03-21 | 59295.00581 | -13.8064 | 0.0021 | 1.48 | 2.1 | 9.73e+03 | 5.57e+03 | 1.10e+07 | 1.11e+07 | 5.40e+04 | 3.09e+04 | -2.5 | 0.4 |
| 2021-03-22 | 59296.06857 | -13.8122 | 0.0022 | -4.33 | 2.2 | -9.15e+03 | 5.67e+03 | 9.05e+06 | 1.11e+07 | -5.07e+04 | 3.14e+04 | -8.0 | 0.4 |
| 2021-03-23 | 59297.00310 | -13.8058 | 0.0020 | 2.05 | 2.0 | -8.88e+03 | 5.28e+03 | -4.55e+06 | 1.05e+07 | -4.93e+04 | 2.93e+04 | -10.8 | 0.4 |
| 2021-03-24 | 59298.05160 | -13.8056 | 0.0021 | 2.25 | 2.1 | -1.73e+04 | 5.34e+03 | -1.49e+06 | 1.06e+07 | -9.58e+04 | 2.96e+04 | -13.0 | 0.4 |
| 2021-03-26 | 59300.02528 | -13.8228 | 0.0027 | -14.98 | 2.7 | 5.65e+04 | 6.99e+03 | 1.69e+07 | 1.38e+07 | 3.13e+05 | 3.88e+04 | -1.0 | 0.5 |
| 2021-03-27 | 59300.80238 | -13.8151 | 0.0021 | -7.23 | 2.1 | 2.17e+03 | 5.32e+03 | -1.35e+07 | 1.05e+07 | 1.20e+04 | 2.95e+04 | -1.5 | 0.4 |
| 2021-03-31 | 59304.97415 | -13.8132 | 0.0022 | -5.37 | 2.2 | 5.04e+04 | 5.61e+03 | -1.02e+07 | 1.11e+07 | 2.80e+05 | 3.11e+04 | 13.9 | 0.4 |
| 2021-04-01 | 59306.02154 | -13.8135 | 0.0020 | -5.66 | 2.0 | 5.31e+04 | 5.26e+03 | 4.51e+06 | 1.04e+07 | 2.95e+05 | 2.92e+04 | 12.4 | 0.4 |
| 2021-04-22 | 59326.95738 | -13.8101 | 0.0022 | -2.22 | 2.2 | -5.94e+03 | 5.77e+03 | 1.04e+07 | 1.15e+07 | -3.29e+04 | 3.20e+04 | -8.4 | 0.5 |
| 2021-04-23 | 59327.87646 | -13.7980 | 0.0021 | 9.87 | 2.1 | 1.02e+04 | 5.47e+03 | 1.25e+07 | 1.08e+07 | 5.65e+04 | 3.03e+04 | -5.2 | 0.4 |
| 2021-04-24 | 59328.96195 | -13.8032 | 0.0022 | 4.67 | 2.2 | 1.46e+04 | 5.62e+03 | 9.72e+06 | 1.12e+07 | 8.11e+04 | 3.12e+04 | -1.2 | 0.4 |
| 2021-04-25 | 59329.91740 | -13.8170 | 0.0027 | -9.16 | 2.7 | 2.76e+04 | 6.87e+03 | -7.31e+05 | 1.36e+07 | 1.53e+05 | 3.81e+04 | 1.0 | 0.5 |
| 2021-04-26 | 59330.94314 | -13.8139 | 0.0021 | -6.08 | 2.1 | 3.09e+04 | 5.54e+03 | 1.29e+07 | 1.10e+07 | 1.71e+05 | 3.07e+04 | 3.1 | 0.4 |
| 2021-04-27 | 59331.96518 | -13.8093 | 0.0022 | -1.44 | 2.2 | 3.94e+04 | 5.64e+03 | 4.88e+06 | 1.12e+07 | 2.19e+05 | 3.13e+04 | 6.3 | 0.5 |
| 2021-04-28 | 59332.96320 | -13.8151 | 0.0021 | -7.28 | 2.1 | 3.64e+04 | 5.46e+03 | 5.39e+06 | 1.09e+07 | 2.02e+05 | 3.03e+04 | 10.7 | 0.4 |
| 2021-04-30 | 59334.93467 | -13.8178 | 0.0022 | -9.97 | 2.2 | 4.75e+04 | 5.66e+03 | -3.03e+05 | 1.13e+07 | 2.64e+05 | 3.14e+04 | 9.0 | 0.4 |
| 2021-05-01 | 59335.94049 | -13.8179 | 0.0021 | -10.04 | 2.1 | 9.51e+03 | 5.43e+03 | -1.66e+06 | 1.08e+07 | 5.27e+04 | 3.01e+04 | 4.4 | 0.4 |
| 2021-05-02 | 59336.93193 | -13.8165 | 0.0022 | -8.69 | 2.2 | -1.63e+04 | 5.80e+03 | -9.77e+06 | 1.15e+07 | -9.02e+04 | 3.22e+04 | -1.6 | 0.5 |
| 2021-05-03 | 59337.91950 | -13.8201 | 0.0023 | -12.27 | 2.3 | -2.79e+04 | 5.89e+03 | 2.21e+07 | 1.17e+07 | -1.55e+05 | 3.27e+04 | -4.7 | 0.5 |
| 2021-06-19 | 59384.77577 | -13.8130 | 0.0020 | -5.19 | 2.0 | 2.47e+04 | 5.29e+03 | 1.23e+07 | 1.05e+07 | 1.37e+05 | 2.93e+04 | 4.4 | 0.4 |
| 2021-06-20 | 59385.77368 | -13.8039 | 0.0023 | 3.97 | 2.3 | 2.73e+04 | 5.87e+03 | 1.14e+07 | 1.16e+07 | 1.52e+05 | 3.26e+04 | 2.0 | 0.5 |
| 2021-06-21 | 59386.76976 | -13.8045 | 0.0022 | 3.39 | 2.2 | 3.71e+04 | 5.67e+03 | 6.02e+06 | 1.12e+07 | 2.06e+05 | 3.15e+04 | 2.9 | 0.4 |
| 2021-06-23 | 59388.75094 | -13.8013 | 0.0023 | 6.51 | 2.3 | 1.76e+04 | 6.10e+03 | -8.91e+06 | 1.21e+07 | 9.77e+04 | 3.38e+04 | 5.4 | 0.5 |
| 2021-06-24 | 59389.74885 | -13.8187 | 0.0022 | -10.81 | 2.2 | 2.81e+04 | 5.80e+03 | -3.87e+06 | 1.15e+07 | 1.56e+05 | 3.21e+04 | 5.8 | 0.5 |
| 2021-06-25 | 59390.75487 | -13.8069 | 0.0021 | 0.91 | 2.1 | 2.93e+04 | 5.54e+03 | -6.53e+06 | 1.10e+07 | 1.63e+05 | 3.07e+04 | 5.2 | 0.4 |
| 2021-06-26 | 59391.74876 | -13.8123 | 0.0021 | -4.41 | 2.1 | 1.64e+04 | 5.44e+03 | -8.85e+05 | 1.08e+07 | 9.08e+04 | 3.02e+04 | 2.1 | 0.4 |
| 2021-06-27 | 59392.75010 | -13.8161 | 0.0021 | -8.28 | 2.1 | 2.90e+04 | 5.60e+03 | 2.84e+07 | 1.11e+07 | 1.61e+05 | 3.10e+04 | 2.3 | 0.4 |
| 2021-06-28 | 59393.75228 | -13.8045 | 0.0023 | 3.34 | 2.3 | -1.01e+04 | 6.02e+03 | 1.33e+07 | 1.20e+07 | -5.60e+04 | 3.34e+04 | -1.3 | 0.5 |
| 2021-06-29 | 59394.74804 | -13.8086 | 0.0023 | -0.73 | 2.3 | 1.56e+04 | 6.12e+03 | 1.18e+07 | 1.22e+07 | 8.64e+04 | 3.40e+04 | 1.8 | 0.5 |
| 2021-06-30 | 59395.74896 | -13.8157 | 0.0026 | -7.87 | 2.6 | 5.13e+04 | 6.65e+03 | -3.57e+06 | 1.31e+07 | 2.84e+05 | 3.69e+04 | 4.5 | 0.5 |
| 2021-07-01 | 59396.74897 | -13.8208 | 0.0023 | -12.99 | 2.3 | 2.80e+04 | 5.89e+03 | 1.83e+07 | 1.16e+07 | 1.55e+05 | 3.26e+04 | 4.9 | 0.5 |
| 2021-07-02 | 59397.74861 | -13.8168 | 0.0028 | -8.96 | 2.8 | 5.21e+04 | 7.39e+03 | 2.70e+07 | 1.47e+07 | 2.89e+05 | 4.10e+04 | 6.5 | 0.6 |
| 2021-07-17 | 59412.74636 | -13.8160 | 0.0020 | -8.15 | 2.0 | 4.31e+04 | 5.30e+03 | 2.42e+07 | 1.05e+07 | 2.39e+05 | 2.94e+04 | 8.6 | 0.4 |
| 2021-07-19 | 59414.73953 | -13.8202 | 0.0022 | -12.31 | 2.2 | 5.55e+04 | 5.78e+03 | 7.46e+06 | 1.15e+07 | 3.08e+05 | 3.21e+04 | 10.3 | 0.5 |
| 2021-11-16 | 59535.13288 | -13.8093 | 0.0020 | -1.47 | 2.0 | 2.65e+04 | 5.18e+03 | 8.20e+06 | 1.03e+07 | 1.47e+05 | 2.87e+04 | 8.1 | 0.4 |
| 2021-11-17 | 59536.15385 | -13.8130 | 0.0021 | -5.19 | 2.1 | 1.60e+04 | 5.53e+03 | -5.77e+05 | 1.10e+07 | 8.89e+04 | 3.07e+04 | 7.1 | 0.4 |
| 2021-11-18 | 59537.12805 | -13.8037 | 0.0020 | 4.14 | 2.0 | 2.26e+04 | 5.23e+03 | 6.64e+06 | 1.04e+07 | 1.25e+05 | 2.90e+04 | 7.1 | 0.4 |
| 2021-11-19 | 59538.15074 | -13.7934 | 0.0022 | 14.46 | 2.2 | 2.02e+04 | 5.70e+03 | -2.70e+07 | 1.13e+07 | 1.12e+05 | 3.16e+04 | 8.0 | 0.4 |
| 2021-11-20 | 59539.15254 | -13.8053 | 0.0015 | 2.55 | 1.5 | 4.60e+04 | 3.77e+03 | 1.87e+06 | 7.44e+06 | 2.55e+05 | 2.09e+04 | 11.0 | 0.3 |
| 2021-11-21 | 59540.13474 | -13.8126 | 0.0019 | -4.78 | 1.9 | 4.74e+04 | 5.02e+03 | 2.60e+06 | 9.90e+06 | 2.63e+05 | 2.78e+04 | 11.0 | 0.4 |
| 2021-11-22 | 59541.12292 | -13.8061 | 0.0021 | 1.79 | 2.1 | 4.04e+04 | 5.55e+03 | 2.70e+06 | 1.09e+07 | 2.24e+05 | 3.08e+04 | 11.4 | 0.4 |
| 2021-12-09 | 59558.14083 | -13.8122 | 0.0020 | -4.36 | 2.0 | 2.17e+04 | 5.20e+03 | 1.42e+07 | 1.03e+07 | 1.20e+05 | 2.89e+04 | 3.3 | 0.4 |
| 2021-12-10 | 59559.12174 | -13.8144 | 0.0023 | -6.56 | 2.3 | 2.94e+03 | 5.98e+03 | 1.03e+07 | 1.19e+07 | 1.63e+04 | 3.32e+04 | 1.0 | 0.5 |
| 2021-12-11 | 59560.16381 | -13.8097 | 0.0023 | -1.83 | 2.3 | 1.52e+04 | 5.94e+03 | 5.63e+06 | 1.18e+07 | 8.42e+04 | 3.29e+04 | -0.4 | 0.5 |
| 2021-12-13 | 59562.12940 | -13.8001 | 0.0021 | 7.80 | 2.1 | 3.96e+04 | 5.37e+03 | -4.65e+06 | 1.07e+07 | 2.19e+05 | 2.98e+04 | 3.8 | 0.4 |
| 2021-12-15 | 59564.11415 | -13.8091 | 0.0020 | -1.29 | 2.0 | 1.27e+04 | 5.21e+03 | 6.83e+06 | 1.03e+07 | 7.03e+04 | 2.89e+04 | 2.9 | 0.4 |
| 2021-12-16 | 59565.07512 | -13.8103 | 0.0022 | -2.48 | 2.2 | 1.46e+04 | 5.81e+03 | -1.39e+07 | 1.15e+07 | 8.11e+04 | 3.22e+04 | 3.5 | 0.5 |
| 2021-12-17 | 59566.16860 | -13.8094 | 0.0029 | -1.59 | 2.9 | 5.43e+04 | 7.65e+03 | 7.99e+06 | 1.52e+07 | 3.01e+05 | 4.24e+04 | 5.5 | 0.6 |
| 2021-12-18 | 59567.13357 | -13.8133 | 0.0022 | -5.40 | 2.2 | 1.93e+04 | 5.68e+03 | 1.13e+07 | 1.13e+07 | 1.07e+05 | 3.15e+04 | 9.4 | 0.4 |
| 2022-01-06 | 59586.13516 | -13.8073 | 0.0020 | 0.60 | 2.0 | 1.92e+04 | 5.17e+03 | 4.33e+06 | 1.03e+07 | 1.06e+05 | 2.87e+04 | 6.0 | 0.4 |
| 2022-01-08 | 59588.15572 | -13.8158 | 0.0019 | -7.98 | 1.9 | 2.46e+04 | 5.05e+03 | -9.02e+06 | 9.99e+06 | 1.36e+05 | 2.80e+04 | 3.2 | 0.4 |
| 2022-01-09 | 59589.10563 | -13.8113 | 0.0021 | -3.40 | 2.1 | 3.00e+04 | 5.47e+03 | -2.34e+06 | 1.08e+07 | 1.66e+05 | 3.03e+04 | 1.7 | 0.4 |
| 2022-01-10 | 59590.13419 | -13.8130 | 0.0022 | -5.11 | 2.2 | 3.39e+04 | 5.90e+03 | -3.40e+06 | 1.17e+07 | 1.88e+05 | 3.27e+04 | 1.7 | 0.5 |
| 2022-01-12 | 59592.18938 | -13.7988 | 0.0020 | 9.05 | 2.0 | 5.05e+04 | 5.27e+03 | 1.82e+07 | 1.05e+07 | 2.80e+05 | 2.92e+04 | -4.0 | 0.4 |
| 2022-01-13 | 59593.11775 | -13.8072 | 0.0020 | 0.69 | 2.0 | -8.79e+03 | 5.20e+03 | 3.61e+06 | 1.03e+07 | -4.87e+04 | 2.89e+04 | -4.2 | 0.4 |
| 2022-01-14 | 59594.10650 | -13.8182 | 0.0019 | -10.35 | 1.9 | 9.62e+03 | 4.95e+03 | 3.38e+06 | 9.80e+06 | 5.33e+04 | 2.74e+04 | 0.0 | 0.4 |
| 2022-01-15 | 59595.18542 | -13.8149 | 0.0020 | -7.06 | 2.0 | 2.78e+04 | 5.20e+03 | 7.46e+06 | 1.03e+07 | 1.54e+05 | 2.89e+04 | 6.9 | 0.4 |
| 2022-01-16 | 59596.12355 | -13.8143 | 0.0026 | -6.44 | 2.6 | 5.14e+04 | 6.73e+03 | 1.51e+07 | 1.33e+07 | 2.85e+05 | 3.73e+04 | 12.2 | 0.5 |
| 2022-01-17 | 59597.13243 | -13.8054 | 0.0020 | 2.49 | 2.0 | 4.43e+04 | 5.13e+03 | 9.68e+06 | 1.02e+07 | 2.46e+05 | 2.85e+04 | 13.2 | 0.4 |
| 2022-01-18 | 59598.17872 | -13.8049 | 0.0020 | 2.97 | 2.0 | 4.90e+04 | 5.22e+03 | 1.15e+07 | 1.03e+07 | 2.72e+05 | 2.89e+04 | 12.0 | 0.4 |
| 2022-01-19 | 59599.18869 | -13.8110 | 0.0020 | -3.16 | 2.0 | 2.96e+04 | 5.23e+03 | 8.70e+06 | 1.04e+07 | 1.64e+05 | 2.90e+04 | 7.7 | 0.4 |
| 2022-01-20 | 59600.18208 | -13.8112 | 0.0021 | -3.33 | 2.1 | 1.92e+04 | 5.55e+03 | 9.47e+06 | 1.10e+07 | 1.06e+05 | 3.08e+04 | 5.4 | 0.4 |
| 2022-01-22 | 59602.13492 | -13.8196 | 0.0026 | -11.76 | 2.6 | 2.17e+04 | 6.65e+03 | 1.18e+07 | 1.31e+07 | 1.20e+05 | 3.69e+04 | 5.1 | 0.5 |
| 2022-01-24 | 59604.13357 | -13.8071 | 0.0020 | 0.72 | 2.0 | 1.89e+04 | 5.35e+03 | 1.36e+07 | 1.06e+07 | 1.05e+05 | 2.96e+04 | 1.8 | 0.4 |
| 2022-01-25 | 59605.15366 | -13.8073 | 0.0020 | 0.59 | 2.0 | 2.51e+04 | 5.12e+03 | 9.71e+05 | 1.01e+07 | 1.39e+05 | 2.84e+04 | -1.2 | 0.4 |
| 2022-01-26 | 59606.10853 | -13.8149 | 0.0019 | -7.05 | 1.9 | 9.48e+03 | 5.04e+03 | 1.53e+07 | 1.00e+07 | 5.26e+04 | 2.80e+04 | -5.8 | 0.4 |
| 2022-03-11 | 59649.87520 | -13.8077 | 0.0020 | 0.14 | 2.0 | -3.36e+04 | 5.22e+03 | -1.98e+07 | 1.04e+07 | -1.86e+05 | 2.89e+04 | -7.8 | 0.4 |
| 2022-03-12 | 59650.96249 | -13.8077 | 0.0021 | 0.11 | 2.1 | 3.64e+03 | 5.36e+03 | -3.71e+04 | 1.07e+07 | 2.02e+04 | 2.97e+04 | -8.3 | 0.4 |
| 2022-03-13 | 59651.98257 | -13.8016 | 0.0021 | 6.24 | 2.1 | 6.33e+03 | 5.53e+03 | -1.49e+07 | 1.10e+07 | 3.51e+04 | 3.07e+04 | -5.7 | 0.4 |
| 2022-03-14 | 59652.86144 | -13.8082 | 0.0021 | -0.31 | 2.1 | 8.88e+03 | 5.52e+03 | 1.40e+07 | 1.09e+07 | 4.93e+04 | 3.06e+04 | -1.6 | 0.4 |
| 2022-03-15 | 59654.02433 | -13.8117 | 0.0021 | -3.88 | 2.1 | 7.32e+03 | 5.44e+03 | -9.61e+06 | 1.08e+07 | 4.06e+04 | 3.02e+04 | 2.2 | 0.4 |
| 2022-03-16 | 59655.06137 | -13.8097 | 0.0021 | -1.83 | 2.1 | 2.70e+04 | 5.39e+03 | -2.58e+06 | 1.07e+07 | 1.50e+05 | 2.99e+04 | 4.3 | 0.4 |
| 2022-03-17 | 59655.97333 | -13.8175 | 0.0022 | -9.67 | 2.2 | 3.02e+04 | 5.62e+03 | -3.20e+06 | 1.12e+07 | 1.67e+05 | 3.12e+04 | 4.2 | 0.4 |
| 2022-03-18 | 59657.06403 | -13.8151 | 0.0025 | -7.20 | 2.5 | 3.03e+04 | 6.39e+03 | 7.81e+06 | 1.27e+07 | 1.68e+05 | 3.54e+04 | 5.3 | 0.5 |
| 2022-03-19 | 59658.05424 | -13.8093 | 0.0021 | -1.49 | 2.1 | 7.23e+03 | 5.35e+03 | 1.74e+06 | 1.06e+07 | 4.01e+04 | 2.97e+04 | 5.2 | 0.4 |
| 2022-03-20 | 59658.99954 | -13.8091 | 0.0020 | -1.29 | 2.0 | 2.72e+04 | 5.25e+03 | -6.57e+06 | 1.04e+07 | 1.51e+05 | 2.91e+04 | 7.3 | 0.4 |
| 2022-03-22 | 59660.99607 | -13.8027 | 0.0021 | 5.13 | 2.1 | 4.06e+04 | 5.47e+03 | 1.18e+06 | 1.08e+07 | 2.25e+05 | 3.03e+04 | 4.0 | 0.4 |
| 2022-03-23 | 59661.99049 | -13.8129 | 0.0017 | -5.07 | 1.7 | 3.23e+04 | 4.37e+03 | 4.23e+06 | 8.70e+06 | 1.79e+05 | 2.43e+04 | -1.0 | 0.3 |
| 2022-04-09 | 59678.99154 | -13.8067 | 0.0021 | 1.14 | 2.1 | 2.87e+04 | 5.57e+03 | -2.13e+07 | 1.11e+07 | 1.59e+05 | 3.09e+04 | -2.8 | 0.4 |
| 2022-04-12 | 59681.97934 | -13.8045 | 0.0021 | 3.33 | 2.1 | 5.24e+04 | 5.52e+03 | 1.74e+07 | 1.10e+07 | 2.91e+05 | 3.06e+04 | 3.7 | 0.4 |
| 2022-04-13 | 59682.99757 | -13.8055 | 0.0021 | 2.40 | 2.1 | 3.27e+04 | 5.36e+03 | 2.90e+07 | 1.07e+07 | 1.81e+05 | 2.97e+04 | 2.6 | 0.4 |
| 2022-04-14 | 59684.01154 | -13.8032 | 0.0030 | 4.70 | 3.0 | 4.75e+04 | 7.71e+03 | 9.58e+06 | 1.54e+07 | 2.63e+05 | 4.28e+04 | 1.8 | 0.6 |
| 2022-04-15 | 59684.98764 | -13.8079 | 0.0022 | 0.00 | 2.2 | 1.62e+04 | 5.71e+03 | 3.79e+05 | 1.14e+07 | 8.99e+04 | 3.17e+04 | 2.3 | 0.5 |
| 2022-04-18 | 59687.98838 | -13.7988 | 0.0022 | 9.03 | 2.2 | 1.02e+04 | 5.76e+03 | 7.79e+06 | 1.15e+07 | 5.68e+04 | 3.19e+04 | -0.5 | 0.5 |
| 2022-04-19 | 59688.96958 | -13.7970 | 0.0021 | 10.84 | 2.1 | 1.76e+04 | 5.60e+03 | 1.91e+07 | 1.12e+07 | 9.77e+04 | 3.11e+04 | 1.3 | 0.4 |
| 2022-04-20 | 59689.95600 | -13.8017 | 0.0025 | 6.15 | 2.5 | -1.17e+04 | 6.60e+03 | 6.51e+06 | 1.32e+07 | -6.46e+04 | 3.66e+04 | -1.7 | 0.5 |
| 2022-04-21 | 59690.95500 | -13.8134 | 0.0025 | -5.57 | 2.5 | 1.21e+04 | 6.57e+03 | -4.24e+07 | 1.31e+07 | 6.73e+04 | 3.64e+04 | -2.4 | 0.5 |
| 2022-05-11 | 59710.90469 | -13.8146 | 0.0022 | -6.79 | 2.2 | 4.60e+04 | 5.66e+03 | 4.43e+06 | 1.12e+07 | 2.55e+05 | 3.14e+04 | 4.7 | 0.4 |
| 2022-05-12 | 59711.91277 | -13.8096 | 0.0021 | -1.70 | 2.1 | 1.51e+04 | 5.50e+03 | 1.32e+07 | 1.09e+07 | 8.40e+04 | 3.05e+04 | 3.6 | 0.4 |
| 2022-05-13 | 59712.88788 | -13.8094 | 0.0022 | -1.59 | 2.2 | 1.60e+04 | 5.66e+03 | 1.16e+07 | 1.12e+07 | 8.86e+04 | 3.14e+04 | 2.2 | 0.4 |
| 2022-05-16 | 59715.85928 | -13.8133 | 0.0021 | -5.45 | 2.1 | -5.31e+03 | 5.59e+03 | 4.11e+06 | 1.11e+07 | -2.95e+04 | 3.10e+04 | -6.6 | 0.4 |
| 2022-05-17 | 59716.86406 | -13.8129 | 0.0024 | -5.01 | 2.4 | 1.33e+04 | 6.20e+03 | -2.64e+07 | 1.22e+07 | 7.38e+04 | 3.44e+04 | -7.1 | 0.5 |
| 2022-06-03 | 59733.77334 | -13.8190 | 0.0022 | -11.17 | 2.2 | 3.58e+04 | 5.81e+03 | -4.83e+06 | 1.16e+07 | 1.99e+05 | 3.22e+04 | -1.8 | 0.5 |
| 2022-06-05 | 59735.77903 | -13.8068 | 0.0021 | 1.08 | 2.1 | 5.98e+04 | 5.53e+03 | -1.28e+07 | 1.10e+07 | 3.31e+05 | 3.07e+04 | 2.4 | 0.4 |
| 2022-06-06 | 59736.84702 | -13.7989 | 0.0021 | 8.94 | 2.1 | 1.27e+04 | 5.53e+03 | -5.40e+06 | 1.10e+07 | 7.02e+04 | 3.07e+04 | 0.6 | 0.4 |
| 2022-06-07 | 59737.85329 | -13.8014 | 0.0021 | 6.41 | 2.1 | 7.48e+03 | 5.55e+03 | -2.09e+06 | 1.10e+07 | 4.15e+04 | 3.08e+04 | -3.0 | 0.4 |
| 2022-06-08 | 59738.75236 | -13.8115 | 0.0021 | -3.65 | 2.1 | 2.03e+04 | 5.41e+03 | -2.12e+04 | 1.06e+07 | 1.12e+05 | 3.00e+04 | -2.7 | 0.4 |
| 2022-06-09 | 59739.75308 | -13.8116 | 0.0022 | -3.79 | 2.2 | 1.13e+04 | 5.65e+03 | -1.61e+07 | 1.10e+07 | 6.27e+04 | 3.13e+04 | -3.9 | 0.4 |
| 2022-06-10 | 59740.79830 | -13.8142 | 0.0022 | -6.33 | 2.2 | -1.49e+04 | 5.62e+03 | 1.70e+07 | 1.10e+07 | -8.28e+04 | 3.12e+04 | -4.1 | 0.4 |
| 2022-06-11 | 59741.78115 | -13.8021 | 0.0023 | 5.74 | 2.3 | 1.94e+04 | 5.97e+03 | 6.14e+06 | 1.18e+07 | 1.08e+05 | 3.31e+04 | -4.9 | 0.5 |
| 2022-06-12 | 59742.79232 | -13.8000 | 0.0027 | 7.90 | 2.7 | -1.77e+04 | 7.01e+03 | -1.27e+07 | 1.36e+07 | -9.79e+04 | 3.89e+04 | -8.0 | 0.5 |
| 2022-06-13 | 59743.75821 | -13.8036 | 0.0031 | 4.20 | 3.1 | 4.26e+01 | 8.09e+03 | 1.48e+07 | 1.59e+07 | 2.36e+02 | 4.48e+04 | -10.3 | 0.6 |
| 2022-11-06 | 59890.13221 | -13.8056 | 0.0022 | 2.24 | 2.2 | -6.02e+03 | 5.59e+03 | 6.75e+05 | 1.11e+07 | -3.34e+04 | 3.10e+04 | -4.0 | 0.4 |
| 2022-11-09 | 59893.12159 | -13.8060 | 0.0019 | 1.84 | 1.9 | 2.27e+04 | 4.96e+03 | -4.66e+06 | 9.73e+06 | 1.26e+05 | 2.75e+04 | -2.0 | 0.4 |
| 2022-11-12 | 59896.11941 | -13.8066 | 0.0022 | 1.27 | 2.2 | -8.55e+03 | 5.60e+03 | -2.28e+07 | 1.11e+07 | -4.74e+04 | 3.10e+04 | -10.6 | 0.4 |
| 2022-11-15 | 59899.15299 | -13.8056 | 0.0020 | 2.27 | 2.0 | 3.63e+04 | 5.15e+03 | 4.48e+05 | 1.02e+07 | 2.01e+05 | 2.86e+04 | -3.3 | 0.4 |
| 2022-11-18 | 59902.15197 | -13.8129 | 0.0021 | -5.07 | 2.1 | 6.36e+03 | 5.51e+03 | -9.23e+06 | 1.09e+07 | 3.52e+04 | 3.05e+04 | -2.4 | 0.4 |
| 2022-11-19 | 59903.15749 | -13.8059 | 0.0020 | 1.95 | 2.0 | 2.32e+04 | 5.24e+03 | -1.45e+07 | 1.04e+07 | 1.29e+05 | 2.91e+04 | -2.4 | 0.4 |
| 2022-12-30 | 59944.16181 | -13.8032 | 0.0022 | 4.62 | 2.2 | 1.33e+04 | 5.85e+03 | 3.62e+06 | 1.16e+07 | 7.36e+04 | 3.24e+04 | -13.4 | 0.4 |
| 2022-12-31 | 59945.15901 | -13.8060 | 0.0022 | 1.87 | 2.2 | -1.08e+03 | 5.59e+03 | -2.06e+07 | 1.10e+07 | -6.00e+03 | 3.10e+04 | -10.7 | 0.4 |
| 2023-01-01 | 59946.15454 | -13.7994 | 0.0021 | 8.42 | 2.1 | 3.26e+04 | 5.40e+03 | -5.34e+06 | 1.06e+07 | 1.81e+05 | 2.99e+04 | -3.3 | 0.4 |
| 2023-01-04 | 59949.14985 | -13.7976 | 0.0022 | 10.23 | 2.2 | 2.03e+04 | 5.75e+03 | -5.11e+06 | 1.13e+07 | 1.12e+05 | 3.19e+04 | 8.9 | 0.4 |
| 2023-01-05 | 59950.17815 | -13.8061 | 0.0021 | 1.80 | 2.1 | 5.57e+04 | 5.32e+03 | -5.70e+06 | 1.05e+07 | 3.09e+05 | 2.95e+04 | 5.3 | 0.4 |
| 2023-01-06 | 59951.14762 | -13.8064 | 0.0019 | 1.49 | 1.9 | 1.91e+04 | 5.00e+03 | -4.90e+06 | 9.89e+06 | 1.06e+05 | 2.77e+04 | -2.0 | 0.4 |
| 2023-01-08 | 59953.18602 | -13.8048 | 0.0019 | 3.06 | 1.9 | 2.68e+04 | 5.03e+03 | 1.42e+06 | 9.92e+06 | 1.48e+05 | 2.79e+04 | -6.3 | 0.4 |
| 2023-01-09 | 59954.15923 | -13.7979 | 0.0019 | 9.94 | 1.9 | 3.29e+04 | 4.98e+03 | 1.22e+06 | 9.79e+06 | 1.82e+05 | 2.76e+04 | -5.6 | 0.4 |
| 2023-01-11 | 59956.18776 | -13.8126 | 0.0020 | -4.74 | 2.0 | 1.61e+04 | 5.23e+03 | 4.14e+06 | 1.03e+07 | 8.90e+04 | 2.90e+04 | -9.5 | 0.4 |
| 2023-01-13 | 59958.12996 | -13.7989 | 0.0021 | 8.96 | 2.1 | 1.27e+04 | 5.33e+03 | -1.91e+07 | 1.05e+07 | 7.03e+04 | 2.96e+04 | -17.6 | 0.4 |
| 2023-02-02 | 59978.09021 | -13.7949 | 0.0019 | 13.00 | 1.9 | 5.75e+04 | 4.80e+03 | -1.59e+07 | 9.52e+06 | 3.19e+05 | 2.66e+04 | 5.2 | 0.4 |
| 2023-02-08 | 59984.11968 | -13.7913 | 0.0022 | 16.53 | 2.2 | 3.66e+04 | 5.58e+03 | -1.30e+07 | 1.10e+07 | 2.03e+05 | 3.09e+04 | -1.6 | 0.4 |
| 2023-03-31 | 60035.00130 | -13.8007 | 0.0019 | 7.11 | 1.9 | 4.06e+04 | 5.05e+03 | 2.49e+07 | 1.00e+07 | 2.25e+05 | 2.80e+04 | -4.0 | 0.4 |
| 2023-04-07 | 60041.95876 | -13.7906 | 0.0020 | 17.23 | 2.0 | 3.52e+04 | 5.24e+03 | 4.43e+06 | 1.04e+07 | 1.95e+05 | 2.91e+04 | -8.7 | 0.4 |
| 2023-04-11 | 60045.99173 | -13.8129 | 0.0021 | -5.04 | 2.1 | 4.71e+04 | 5.38e+03 | 1.97e+07 | 1.07e+07 | 2.61e+05 | 2.98e+04 | -5.3 | 0.4 |
| 2023-04-29 | 60063.90176 | -13.8025 | 0.0020 | 5.34 | 2.0 | 7.43e+04 | 5.17e+03 | -1.60e+06 | 1.02e+07 | 4.12e+05 | 2.87e+04 | -3.7 | 0.4 |
| 2023-05-02 | 60066.90070 | -13.8122 | 0.0024 | -4.35 | 2.4 | 2.34e+04 | 6.38e+03 | 1.01e+05 | 1.26e+07 | 1.30e+05 | 3.54e+04 | -9.2 | 0.5 |
| 2023-05-05 | 60069.86886 | -13.8057 | 0.0022 | 2.12 | 2.2 | 2.99e+04 | 5.79e+03 | -2.98e+07 | 1.15e+07 | 1.66e+05 | 3.21e+04 | -9.8 | 0.5 |
| 2023-05-09 | 60073.87663 | -13.8140 | 0.0022 | -6.10 | 2.2 | 4.97e+04 | 5.81e+03 | -5.37e+06 | 1.15e+07 | 2.76e+05 | 3.22e+04 | -2.0 | 0.5 |
| 2023-06-03 | 60098.77913 | -13.7937 | 0.0022 | 14.12 | 2.2 | 1.56e+04 | 5.68e+03 | -2.07e+07 | 1.13e+07 | 8.63e+04 | 3.15e+04 | -9.2 | 0.4 |
| 2023-06-07 | 60102.79676 | -13.8146 | 0.0025 | -6.77 | 2.5 | 3.87e+04 | 6.46e+03 | -1.38e+07 | 1.28e+07 | 2.15e+05 | 3.58e+04 | 2.8 | 0.5 |
| 2023-06-28 | 60123.74194 | -13.8068 | 0.0023 | 1.01 | 2.3 | 3.24e+04 | 5.93e+03 | 2.28e+07 | 1.17e+07 | 1.80e+05 | 3.29e+04 | -17.2 | 0.5 |
| 2023-07-01 | 60126.73580 | -13.8054 | 0.0020 | 2.49 | 2.0 | 2.53e+04 | 5.27e+03 | 1.57e+06 | 1.04e+07 | 1.40e+05 | 2.92e+04 | -12.4 | 0.4 |
| 2023-07-04 | 60129.75396 | -13.8188 | 0.0029 | -10.94 | 2.9 | 8.80e+04 | 7.55e+03 | 2.29e+07 | 1.50e+07 | 4.88e+05 | 4.19e+04 | 4.4 | 0.6 |
| 2023-07-17 | 60142.73530 | -13.8260 | 0.0023 | -18.10 | 2.3 | 6.82e+04 | 6.01e+03 | 5.14e+07 | 1.19e+07 | 3.78e+05 | 3.33e+04 | -1.5 | 0.5 |
Appendix C Activity Analysis
We measured activity indicators in the SOPHIE and SPIRou spectra. For the SOPHIE data-set, the activity indicators were derived based on distortions of the CCF (bisector (BIS), the Full-Width-Half Maximum (FWHM) and the CCF contrast), and on the spectral line indexes H and S index (based on the Ca II H&K lines, calculated as described by Boisse et al. 2009). We investigated the temporal evolution of the activity indicators calculating GLS periodograms and searching for modulations related to the stellar rotation. The periodograms of the SOPHIE activity indicators are shown in Fig. 18 and 19. Considering the whole ensemble of SOPHIE activity indicators, none of their periodograms display a peak at the stellar rotation period (13.91 days), with exception of Hα and the S index that show excess power below 0.1% FAP around 15 and 16 days, periods which are relatively close to the stellar rotation period. To check for consistency, we also computed GLS periodograms for the SOPHIE data-set split in two periods: 20102012 and 20212023. The periodograms of the 20102012 period do not display significant periodicities except for H which exhibits several peaks at periods between 14 and 19 days, however, they are most likely be due to poor sampling. In the case of the 20212023 period, the periodogram of the S index exhibits a significant periodicity (¡1 FAP) at 15.3 days and more importantly, the BIS shows excess power almost at 1 FAP at 14 days, which is consistent with the rotation period.
For the SPIRou data-set, we analyzed the GLS periodograms of the differential line width (dLW), the FWHM, and the chromatic velocity slope calculated by the LBL algorithm. The periodograms are shown in Fig. 20. We found excess power at 13 days in the dLW and chromatic velocity slope periodograms. This periodic variability could be related to the stellar rotation. No periodicities are detected at 6.02, 18.7 nor 2.99 days neither in the SOPHIE nor SPIRou data-sets.
Appendix D GP modeling of SPIRou’s longitudinal magnetic field.
Measurements of the longitudinal magnetic field () of Gl 410 taken with SPIRou have been previously presented and analyzed in Fouqué et al. (2023) and Donati et al. (2023). Fig. 21 displays the LS periodogram of the data from Donati et al. (2023). We retrieve the peak at days which is associated with to 13.910.09 day rotational period of the star and its 1yr alias at 14.98 days. One additional peak with FAP level below 0.1%, and not related to the window function, is detected at 9.41 days with its 1yr alias at 9.69 days. No peak is observed at 6 days. Using the python module george (Foreman-Mackey 2015), we fitted a GP on the SPIRou time-series using a quasi-periodic kernel of the form:
where is the amplitude, is the smoothing parameter, is the decay time in days, is the period (i.e., the cycle length) in days, and is the uncorrelated noise (s in the corner plots). We used uniform distributions for the priors of the amplitude, decay time, smoothing, cycle length, and uncorrelated noise as described in Table 13. The distribution of the hyperparameters was sampled using the Markov chain Monte Carlo (MCMC) implemented in the EMCEE tool (Foreman-Mackey et al. 2013). We used 100000 samples and 10000 burn-in samples. In Table 13 and Fig. 21, we summarize the results of the GP modeling of the time-series. Fig. 22 present the corner plots of the fit. As expected, we retrieve in our GP the stellar rotational period of days which is consistent with the previous determinations of the rotational period by Donati et al. (2023) and Fouqué et al. (2023). The values obtained for the amplitude, decay time and smoothing parameter are consistent to those of Donati et al. (2023).
| Parameter | Prior distribution | Posterior distribution | ||
| GP + Keplerian RadVel model of the SOPHIE RVs (circular orbit) | ||||
| GP | ||||
| Offset [m s-1] | ||||
| Amplitude [m s-1] | ||||
| Decay time [days] | ||||
| Cycle length (period) [days] | Gaussian | |||
| Smoothing parameter | ||||
| Uncorr. Noise [m s-1] | ||||
| Keplerian | ||||
| Period [days] | ||||
| Semi-amplitude [m s-1 ] | ||||
| Time of inferior conjunction [days] | ||||
| Time of periastron [days] | ||||
| Semi-major axis [au] | ||||
| [M⊕] | ||||
| GP of SPIRou’s measurements | ||||
| ln units | linear units | |||
| Amplitude [G] | ||||
| Decay time [days] | ||||
| Smoothing parameter | ||||
| Cycle length (period) [days] | ||||
| Uncorr. Noise [G] | ||||
| GP of the SOPHIE optical RVs using SPIRou’s as activity proxy | ||||
| Offset [m s-1] | median(RV) + | |||
| Amplitude [m s-1] | ||||
| Decay time [days] | Posterior | |||
| Smoothing parameter | Posterior | |||
| Cycle length (period) [days] | Posterior | |||
| Uncorr. Noise [m s-1] | ||||
Appendix E Corner plots and additional figures and tables of the one-planet model.
Appendix F Corner plots and additional figures and tables of the RadVel 2-planet model.
| Parameter | Prior | Initial guess |
| Gl 410b | ||
| [days] | 6.02 | |
| [m s-1 ] | 4.4 | |
| [JD] | 59648.0 | |
| 0.9 | ||
| 0.0 | ||
| candidate signal at 18 days | ||
| [days] | 18.8† | |
| [m s-1 ] | 3.8 | |
| [JD] | 59652.0 | |
| 0.1 | ||
| 0.0 | ||
| 0.0 |
| rms | BIC | AICc | AICc | ||
| SPIRou wPCA RVs | |||||
| AICc Favored Model | |||||
| , , , , | 10 | 5.3 | 1018 | 988.5 | 0.0 |
| Nearly Indistinguishable | |||||
| , , , | 8 | 5.4 | 1013 | 989.8 | 1.3 |
| Somewhat Disfavored | |||||
| , , , , , | 12 | 5.2 | 1026 | 991.6 | 3.0 |
| , , , , | 10 | 5.3 | 1021 | 992.4 | 3.8 |
| Ruled Out | |||||
| , | 2 | 6.8 | 2422 | 2416 | 1427 |
| , , , | 7 | 20 | 2447 | 2426 | 1438 |
| , , , | 7 | 49 | 2448 | 2427 | 1439 |
| , , | 5 | 346 | 2482 | 2467 | 1479 |
| , , | 5 | 352 | 2583 | 2469 | 1480 |
| SPIRou wPCA | Gl 410b | candidate signal at 18.7 days |
| Orbital parameters | ||
| [days] | ||
| [m s-1] | ||
| [JD] | ||
| Planet parameters | ||
| [M⊕] | ||
| [au] |
Appendix G Dynamical stability analysis
While somehow limited constraints on the orbital configuration can be obtained from the radial velocity measurements, we performed a dynamical stability assessment for the two-planet system Gl 410b and 18.7 day candidate to consolidate the picture on the system architecture. The period ratio (larger than 3) is wide and Gl 410 b and the 18.7d day candidate are small enough that the system is safely stable for circular orbits and the nominal eccentricity values. Indeed the orbital separation in Hill radii is of which is far beyond the stability limit of (Gladman 1993). Although we favor a model with circular orbits, we can derive some constraints on the maximum stable eccentricities by computing the general Hill stability criterion as described in Petit et al. (2018). We take the RadVel posterior and compute the Angular Momentum Deficit (AMD) of the system as well as the critical value ensuring Hill stability (in other words, the maximum AMD such that close encounters are always forbidden). In Fig. 30, we compare the posterior eccentricity distribution resulting from the RadVel RV fit and the posterior eccentricity distribution only keeping the stable configurations. Rejecting Hill unstable configurations from the posterior, we obtain an upper limit for at 0.4 but no additional constraints on . Stable configurations are on average more circular. A similar analysis on the masses show that no stringer constraints can be drawn from the stability on this system.
The Hill stability can also constrain the maximum mutual inclination. To obtain this constraint, we assume that the system invariant plan is aligned with the line of sight191919As the main constraint comes from the mutual inclination and not the impact on the real planetary masses, this is a reasonable hypothesis. and we assign a random mutual inclination to each configuration in the posterior. Computing the critical AMD in this case, we find that systems with a mutual inclination smaller than 40 degrees are stable. As for the eccentricities, this is mainly an upper limit. The system is fully consistent with coplanar orbits.
Appendix H Additional RV analysis of the SPIRou wPCA data.
H.1 periodogram analysis
The periodogram has several inputs: the RV time series to be analysed, its nominal uncertainties, a grid of frequencies , a linear model assumed to be in the data , an assumed covariance matrix of the noise as input. Overall, the model of the RV time series can be described by the equation
where is a Gaussian noise with a certain covariance. Overall, the free parameters of the models are the coefficients of the linear model, , the coefficients of periodic signals , and the covariance parameters, which describe the properties of the noise. In practice, one chooses a parametrization of the covariance matrix of the noise , where the element of at index depends on and a vector of parameters . In the following analysis, the parametrization chosen for is such that its element at index is
where is the nominal measurement uncertainty, is an additional white noise, is a calibration noise, and where equals one if measurements and are taken within the same night and zero otherwise. The quantities and parametrize a correlated noise, which might originate from the star or the instrument. Additionally, , and parametrize a quasi-periodic covariance, potentially resulting from spots or faculae.
The periodogram looks for a small number of frequencies such that the model in Eq. (3) approximates the data. Depending on the different assumptions on the noise model and base linear model, the periodogram might find signals at different frequencies. To test the robustness of the results to a change in the model, we adopt the methodology of Hara et al. (2020). We consider a grid of potential noise models. For each of them, we compute the periodogram, and keep the signals that have a false alarm probability below 0.1. For each noise model in the grid, we then have candidate periodic signals. To each couple noise model/periodic signals, we attribute a cross-validation score. The model is fitted on 60% of the data points (the training set), and the likelihood of its prediction on the remaining 40% (the test set) is computed. This operation is repeated for 300 random draws of selection of training sets, and the final cross validation score is computed as the average of the likelihood of the test set. We also rank the model according to their Bayesian evidence, evaluated with the Laplace approximation (see Hara et al. 2020; Nelson et al. 2020).
In our analysis, we adopt as an eight-column matrix. Its columns are: an offset, the spectral shape indicators (i.e., stellar activity proxies): the SPIRou longitudinal magnetic field (), and the LBL-derived (Artigau et al. 2022, 2024) projection onto the 2nd derivative of the spectrum (d2v), projection onto the 3rd derivative of the spectrum (d3v), differential line width (dLW), and differential temperature with respect to a T=4000 K template (dTEMP4000) and, finally, a cosine and sine term at the sidereal day (period equal to 0.997 days). The grid of values chosen for the noise model is given in Table 19. All the possible combinations of the values of its components were generated, and the corresponding covariance matrices were created according to Eq. (4). In 31, we represent the histogram of cross validation scores.
The periodogram corresponding to the highest ranked model is represented in Fig. 11. It shows four clear peaks at 6, 3, 18 and 1.3 days. To evaluate the robustness of these detections to the noise model, for each the 20% highest ranked models, we report the false alarm probabilities of the peaks found. The false alarms of the highest ranking models are reported in Fig. 32. The percentage of cases where the 6, 3, 18 and 1.3 days signals were found in the 20% best models, the FAPs corresponding to the highest ranked model, and the median FAPs are reported in Table LABEL:tab:faps_cv. We find that the 6, 3, 18 days signal are present in a 100% of the highest ranked models, with median false alarm, probabilities of , and . The highest ranked models by Bayesian evidence give a similar picture, although they seem to provide more statistical significance for the signals (see Table 18). They can be considered respectively as a confirmed planet, a very strong candidate and a strong candidate.
| Period (d) | FAP (best fit) | Inclusion in the model | median FAP |
| 2.997 | 100.0% | ||
| 6.021 | 100.0% | ||
| 18.70 | 100.0% |
| Period (d) | FAP (best fit) | Inclusion in the model | median FAP |
| 2.997 | 100.0% | ||
| 6.021 | 100.0% | ||
| 18.70 | 100.0% |
| Parameter | Values | Highest CV score | Highest evidence |
| (m/s) | 1, 1.5, 2, 2.5, 3, 3.5, 4 | 2.5 | 2 |
| (m/s) | 0, 1 ,1.5, 2 ,2.5, 3, 3.5, 4 | 3 | 2.5 |
| (d) | 0, 3, 6 | 6 | 6 |
| (m/s) | 0, 1 ,1.5, 2 ,2.5, 3, 3.5, 4 | 3 | 2.5 |
| (d) | 13.91, 27.82, 41.73, 55.64 | 13.91 | 27.82 |
| 13.91 | 13.91 | 13.91 | 13.91 |
H.2 Apodized sine periodogram analysis.
To further examine the properties of the signals, we apply the method of Hara et al. (2022), which aim at testing whether the signals have a consistent amplitude, phase and frequency over time, which are the properties expected for planetary signals. For this we compute apodized sine periodograms (ASPs), defined as follows.
We compare the of a linear base model , and the of a model defined as the linear model of plus an apodized sinusoid . The ASP is defined as the difference between the two models
We use the same base model as in our analysis with the periodogram, and use the covariance which corresponds to the noise model with highest Bayesian evidence.
To compute the ASP, we consider a grid of timescales and , where is the total observation timespan of the data. In Fig. 34.a, left panel, we represent as defined in Eq. (5), maximised over for the values of in the grid. In other words, we always select the center of the time window which best fits the data.
We search iteratively for signals. Once the maximum peak of the ASP is found, then we include the corresponding model in the base linear model , and repeat the process. The four first iterations are shown in Fig. 34, a, b, c ,d respectively. The left panels represent the periodograms, the middle panels a zoom in on the highest peak, and the right hand panel represents a statistical significance test, explained below.
In Fig. 34.a, the dominating peak has a period of 6.02 days. We want to test whether the signal is statistically compatible with a constant one (i.e. with ). Denoting by the value of maximising the value of the periodogram (5) for a given and , we compute the distribution of
with the hypothesis that the model is correct, where the fitted cosine and sine amplitudes are obtained by fitting model to the data. can easily be expressed as a generalised distribution, with mean and variance is given by an analytical expression (Hara et al. 2022).
We find that even if the best fitting model for the 6.02 signal is not the purely periodic signal, it is still compatible with a purely periodic one. In Fig. 33.b and c we see that the candidate signals at 2.99 and 18 days are compatible with strictly sinusoidal signals. In the case of the former, other signal shapes are excluded.
In the next iteration, in Fig. 33.d, we find that the best fitting model is a short term signal at 5.78 days. In Fig. 33, we show the corresponding model, which points to a transient signal. possibly caused by outliers or stellar variability.
Since the d3v vector, used in our linear base model , has shown in other SPIRou dataset to be less reliable202020https://lbl.exoplanets.ca/output-data we re-did the whole analysis with the and ASP periodograms without it. The only notable difference is that it seems to slightly decrease the stability of the 18.7 day signal in frequency, amplitude and phase. Let us note that the 2.99 day planet could be in 2:1 mean motion resonance with the 6 d planet. In Hara et al. (2019), it has been shown that eccentric planets can mimic 2:1 mean motion resonances. However, it is very unlikely the case here, as the 2.99 d signal is significant and sharp. In conclusion, both 18.7 and 2.99 d planets are viable candidates, but the evidence for the latter is stronger.